### v1

# LoVR: A Benchmark for Long Video Retrieval in Multimodal Contexts

Abstract
Long videos contain a vast amount of information, making video-text retrieval an essential and challenging task in multimodal learning. However, existing benchmarks suffer from limited video duration, low-quality captions, and coarse annotation granularity, which hinder the evaluation of advanced video-text retrieval methods. To address these limitations, we introduce LoVR, a benchmark specifically designed for long video-text retrieval. LoVR contains 467 long videos and over 40,804 fine-grained clips with high-quality captions. To overcome the issue of poor machine-generated annotations, we propose an efficient caption generation framework that integrates VLM automatic generation, caption quality scoring, and dynamic refinement. This pipeline improves annotation accuracy while maintaining scalability. Furthermore, we introduce a semantic fusion method to generate coherent full-video captions without losing important contextual information. Our benchmark introduces longer videos, more detailed captions, and a larger-scale dataset, presenting new challenges for video understanding and retrieval. Extensive experiments on various advanced embedding models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research. We release the code and dataset link at https://github.com/TechNomad-ds/LoVR-benchmark
Figure 1: Example of video data from LoVR. A long video consists of multiple clips, three of which are shown on the left. The corresponding caption is displayed on the right, describing each clip’s scene and atmosphere, their inter-linkage, and an overall theme summary.
Table 1: Comparison of statistics of video-text benchmarks.
Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator Theme Info.
MSR-VTT xu2016msr 1,000 15.01 9.41 Human
DiDeMo anne2017localizing 1,037 53.94 29.11 Human
MSVD chen2011collecting 670 10.04 7.01 Human
ActivityNet yu2019activitynet 5,044 36.00 13.48 Human
DREAM-1K wang2024tarsier 1,000 8.90 59.30 Human
VDC chai2024auroracap 1,000 28.18 500.91 VLM
CAREBENCH xu2024fine 1,000 14.35 227.95 Human
LoVR
Clip-level 40,804 17.86 1393.65 VLM + Human $\checkmark$
Video-level 467 1560.30 106075.39 VLM + Human $\checkmark$
1 Introduction
In the data-driven era, with the rapid advancement of digital technology, video has become a crucial medium for information transmission due to its intuitive and richly expressive nature pouyanfar2018multimedia . As a core task in multimodal understanding, video-text retrieval, which aims to retrieve semantically relevant video content from large-scale datasets given a textual query, has become increasingly important zhu2023deep .
Recently, videos have grown significantly in length, making retrieving full long-form videos and fine-grained clips increasingly relevant lin2022eclipse ; luo2024video ; eltahir2025multimodal . Consequently, there is an urgent need to benchmark both long video retrieval and clip-level retrieval from long videos to support real-world applications.
However, current video-text retrieval benchmarks are primarily constrained by short video lengths and low-quality captions. These limitations can be summarized in the following three aspects:
(1) Limited Video Length Most widely used datasets—such as MSR-VTT xu2016msr , MSVD chen2011collecting , and DiDeMo anne2017localizing contain only short videos, with average durations less than 1 minute. These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
(2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance. While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) malik2021automatic ), these methods often suffer from low accuracy and semantic ambiguity. Although manual annotation wu2024longvideobench ; song2024moviechat ; yu2019activitynet can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets. Balancing scalability with annotation quality remains a critical challenge in current research.
(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos eltahir2025multimodal ; zhu2024video . This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content. In contrast, constructing clip-level benchmarks by segmenting long videos introduces a more realistic and challenging setting—models must not only retrieve relevant content but also distinguish fine-grained clips from densely packed temporal contexts. This setting better reflects practical applications and poses a greater challenge for retrieval models.
To address the aforementioned limitations, we introduce LoVR, a new benchmark specifically designed for long video-text retrieval. Unlike existing datasets that predominantly consist of short video content, LoVR targets the challenges inherent to long-form videos and aims to support both full-video and fine-grained clip-level retrieval. As shown in Figure 1, the caption of the full video not only contains holistic information but also corresponds to fine-grained clip-level details.
To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency. Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) zhang2024vision , (2) automatically assessing caption quality via the EVQAScore model liang2024evqascore , and (3) conducting a final round of human verification. This hybrid process effectively ensures both high fidelity and scalability of annotations. For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model wang2024qwen2 ; bai2025qwen2 , which demonstrates strong performance on multimodal understanding tasks. The integration of EVQAScore liang2024evqascore further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments. By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios. First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content ge2024v2pe ; zhao2023mmicl . As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
Second, captions play a vital role in aligning textual and visual modalities. In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment chen2025expertized . Accurately modeling high-level semantics—such as overarching themes, narrative progression, and emotional tone—remains an open challenge faghihi2024spectrum .
These observations underscore the importance of a dedicated benchmark. By explicitly targeting long-form content, LoVR surfaces technical bottlenecks in current approaches and provides a valuable foundation for developing more robust long video-text retrieval systems.
Our contributions are as follows:
  1. 1.
    We introduce LoVR, a novel benchmark designed for long video-text retrieval. It comprises 467 full-length videos tailored for long video retrieval, along with 40,804 fine-grained clips to support clip-level retrieval within long videos. Each video and clip is paired with high-quality captions that have undergone both human and machine validation. To the best of our knowledge, LoVR is the first large-scale dataset dedicated to advancing research in long-form video retrieval. Table 1 compares LoVR and other existing benchmarks.
  2. 2.
    To ensure caption quality, we propose a general and cost-effective caption generation pipeline that integrates VLMs with both automated and human validation. Compared to fully automatic or fully manual annotation, our approach achieves high-quality captions while significantly reducing human labor by effectively balancing machine efficiency and human oversight.
  3. 3.
    We conduct systematic experiments using various embedding models on LoVR, providing insights into effective representation learning and retrieval strategies for long-form video-text understanding. Our results reveal that LoVR is more challenging than existing benchmarks, especially in the context of long video retrieval, thus introducing a new challenge and setting a fresh research direction for the community.
Table 2: Duration distribution of long videos in LoVR. ‘Avg.’ refers to the average duration in seconds, while ‘#Clip’ indicates the average number of clips obtained by segmenting each video.
Duration (900, 1000] (1000, 2000] (2000, 3000] (3000, 3600]
Statistics Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip
998.7 4 95.5 1293.2 377 88.6 2569.2 61 84.6 3215.6 25 75.1
2 Related Works
Video-text retrieval benchmarks
Traditional benchmarks (e.g., MSR-VTT xu2016msr , MSVD chen2011collecting , DiDeMo anne2017localizing , TGIF tgif-cvpr2016 , LSMDC lsmdc ) focus on short video clips with concise captions, limiting their ability to evaluate fine-grained semantic understanding. Recent efforts (e.g., ActivityNet yu2019activitynet , HowTo100M miech2019howto100m ) scale up video duration and annotation richness, but HowTo100M relies on noisy ASR-generated captions. To improve annotation quality, DREAM-1K wang2024tarsier and VDC chai2024auroracap use human-annotated descriptions, yet their manual annotation process hinders scalable expansion. To overcome these challenges, we leverage recent advances in multi-concept VLMs an2024mc and visual prompting lin2025drawandunderstand to introduce a VLM-based initial annotator, enabling scalable and fine-grained long-video retrieval.
Video-text retrieval methods
Recent advances have been driven by deep learning techniques. For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations. zhu2023deeplearningvideotextretrieval ; an2024mc Feature matching has evolved from global(e.g. CLIP4Clip luo2021clip4clipempiricalstudyclip ,VSE++ han2022bicnetlearningefficientspatiotemporal ) to multi-granular approaches (local and individual-level alignment)(e.g. CLIP2TV gao2022clip2tvalignmatchdistill ), improving retrieval accuracy. However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
3 LoVR Datasets Construction
The LoVR dataset was constructed through a rigorous process involving raw video selection(Section 3.1) and dataset construction pipeline(Section 3.2). Section 3.3 provides an overview of LoVR and highlights its key advantages in comparison to existing benchmarks.
Figure 2: Overview of the three-stage pipeline in LoVR. Step 1 segments long videos into high-dynamic clips using visual change detection and threshold-based filtering. Step 2 generates high-quality clip-level captions via iterative VLM captioning and human fallback based on EVQAScore. Step 3 constructs long-video captions by clustering clip captions and summarizing them into a full-length description. Finally, a human review process is conducted to ensure quality.
3.1 Video Source
The LoVR benchmark is constructed based on the LongVideoBench wu2024longvideobench dataset. LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 3600 seconds (i.e., 1 hour). The domains of the videos can be broadly categorized into four domains: Life, Movie, Knowledge, and News. To focus on long-form video retrieval, we apply a duration filter and retain only those videos exceeding 15 minutes (900 seconds) in length. Table 2 summarizes the distribution of long videos across different duration intervals, including the average duration and the number of videos in each group.
3.2 Dataset Construction Pipeline
As illustrated in Figure 2, the dataset construction pipeline proposed in this paper consists of three key stages: (1) clip segmentation, (2) caption generation, and (3) full video captioning. This pipeline is designed to enhance both the scalability and the overall quality of the benchmark. Following these stages, final human quality assessment is conducted to ensure the reliability and quality of the benchmark.
3.2.1 Clip Segmentation
Clips serve as fundamental semantic units of videos. Each long video $v\in\mathcal{V}$ is composed of a sequence of temporally contiguous clips, denoted as $v=\{c_{1},c_{2},\dots,c_{N}\}$, where $c_{i}\in\mathcal{C}$. Here, $\mathcal{V}$ denotes the set of all long videos, and $\mathcal{C}$ represents the collection of all clips. Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval. We employ the ContentDetector module from PySceneDetect PySceneDetect , which segments videos based on visual content changes across frames. This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated. We empirically set $\tau=34$, following standard practices in scene detection literature.
Based on visual scene dynamics, videos can be naturally categorized into two types: those with high scene dynamics and those with low scene dynamics. As illustrated in Figure 3, videos such as lectures, interviews, or static photography exhibit minimal visual variation throughout their duration, with scenes remaining largely unchanged across different segments. The inclusion of such videos simplifies the retrieval task, as only a small number of frames may suffice to represent the entire content. In contrast, videos with high scene dynamics contain rich and diverse visual transitions, requiring more precise matching and understanding, which is precisely the level of challenge our benchmark aims to introduce. Furthermore, as shown in the lower portion of Figure 3, the inter-frame difference scores of two representative videos are computed and visualized. The high-scene-dynamics video shows large frame-wise variations, while the low-dynamic one maintains consistently high similarity across frames, often below the threshold for clip segmentation.
To ensure sufficient visual dynamics of the videos, we apply a filtering criterion based on the number of clips generated after segmentation. Specifically, videos yielding fewer than 40 clips are excluded, as this indicates insufficient visual variation in certain segments. These filtered videos often contain numerous long-duration clips with little to no change in content, thereby reducing overall scene dynamics. Finally, we conduct a manual inspection to evaluate the efficacy of our thresholding strategy and confirm its validity. Ultimately, we curated a final set of 467 high-quality long videos that exhibit rich scene dynamics and semantic content diversity, from which a total of 40,804 clips were extracted.
Figure 3: Illustration of Video Dynamics: High vs. Low Scene Dynamics. Example video frames are shown in the upper part (high-dynamic on the left, low-dynamic on the right), while the lower part presents the computed inter-frame difference curves. The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
3.2.2 Caption Generation
We propose a three-stage annotation framework: (1) automatic captioning, (2) quality assessment, and (3) dynamic optimization. This approach integrates VLM-based automatic captioning with human verification, which enables scalable yet precise annotation of clip-level contents.
Automatic Captioning.
We utilize advanced VLMs to produce initial captions based on the visual content of each clip. Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM. To ensure high caption quality, we employ Qwen2.5-VL-Instruct bai2025qwen2 , a series of advanced VLMs. When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style. Here, we emphasize the importance of VLMs in describing theme information, such as overall visual style and atmosphere, which is crucial for understanding videos but often overlooked. As shown in Table 1, existing benchmark captions lack this type of thematic content.
Quality Assessment.
To assess the quality of generated captions, we utilize EVQAScore liang2024evqascore , a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions. For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption. After scoring all instances, we construct a dataset $K=\{(c_{i},s_{i},\text{cap}_{i})\}_{i=1}^{n}$, containing the video clip, its associated caption, and the corresponding quality score.
However, there exists a discrepancy between the computed $s_{i}$ and whether the caption meets human-annotated quality standards. To bridge this gap, we determine an empirical threshold $\theta$ that serves as a benchmark for identifying captions that satisfy manual verification criteria. Specifically, we sample a subset $K^{\prime}\subseteq K$ and conduct manual annotation to count captioning errors per instance. This allows us to establish an empirical correlation between the computed matching scores and human-verified relevance. As illustrated in the left portion of Figure 4, a significant rise in annotation errors is observed for samples with scores below 0.2. Based on this analysis, we set the quality threshold to $\theta=0.2$. Captions with scores above $\theta$ are retained as high-quality, while those below are flagged as low-quality, needing refinement or exclusion.
Figure 4: Left: Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$. The score distribution refers to the left y-axis, while the Content Error Median and Theme Error Median refer to the right y-axis. A significant drop in annotation errors is observed when the match score exceeds 0.2. Middle: Caption length distribution of the long videos. Video-level captions are very long, centered around 10,000 tokens. Right: Caption length distribution of the clips.
Dynamic Optimization.
For captions identified as low-quality, we switch to an alternative VLM to regenerate captions, which are then re-evaluated using the EVQAScore liang2024evqascore . Only those captions achieving a score above the quality threshold $\theta$ are retained. This strategy leverages the diversity among VLMs—arising from differences in training data and architectural design—to improve caption quality for challenging clips. It is only when the number of failed attempts exceeds three that we resort to manual annotation to generate the final captions. This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
3.2.3 Full Video Captioning
Due to the length limitations of VLMs, directly generating captions for entire long videos often leads to incomplete or fragmented descriptions. To address this issue, we sequentially merge adjacent captions by aligning and fusing semantically related segments rather than naively concatenating all clip-level captions or processing them in an LLM beyond its sequence limit. The process is shown as follows:
$\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$ (1)
where LLM denotes the rewriting process of the LLM. In our strategy, we connect the last paragraph of one clip with the first paragraph of the subsequent clip through a language model. During this process, the clip-level captions $\text{cap}_{c}$ are sequentially aggregated to construct the final video-level caption $\text{cap}_{v}$.
3.2.4 Final Human Quality Assessment
To assess the quality of the annotations in LoVR, we conducted a human evaluation study involving 25 participants. A total of 300 clips and 100 long videos were randomly sampled for evaluation. Each sample consisted of a video (or clip) and its associated caption. Participants were instructed to rate the relevance between the visual content and the caption on a 0–5 scale, where a score of 5 denotes perfect semantic alignment with no factual or descriptive errors.
The evaluation results indicate that the captions are of consistently high quality: the average rating across all samples was 4.3, and over 78% of the samples received scores of 4 or 5. These findings validate the effectiveness of our annotation pipeline and demonstrate that the captions in LoVR accurately reflect the underlying video content.
Table 3: Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
Model Text-to-Video Retrieval Text-to-Clip Retrieval
R@1 R@5 R@10 Time w/ IO Time w/o IO R@1 R@5 R@10 Time w/ IO Time w/o IO
Image Encoder
CLIP 69.0 83.0 89.1 21,877,283 3,709,556 22.4 40.6 51.0 22,494,235 3,814,168
CLIP(avg. feat.) 46.0 79.0 89.0 21,877,283 3,709,556 25.7 51.6 61.4 22,494,235 3,814,168
BLIP 39.2 59.0 75.0 174,762 41,746 24.7 47.1 55.7 2,595,565 607,738
ResNet 36.3 58.0 63.0 21,096,459 3,016,545 15.7 27.4 36.0 21,691,391 3,101,613
ResNet(avg. feat.) 31.0 72.0 82.0 21,096,459 3,016,545 15.1 33.9 44.3 21,691,391 3,101,613
Video Encoder
VideoClip 47.2 74.3 82.1 481,926 6,379 45.7 68.2 81.6 41,867,568 529,587
VideoClip-XL 53.0 80.0 86.1 491,180 6,557 52.5 75.6 84.4 42,835,824 571,837
VideoClip-XL-v2 57.0 81.2 87.0 491,190 6,570 56.5 79.2 85.4 42,836,702 573,026
Multimodal Embedding Model
GME-Qwen2-VL-7B 61.2 73.0 79.0 105,823,692 91,234,289 51.2 68.7 78.4 104,221,391 89,229,158
MM-Embed 58.1 71.2 78.0 103,957,148 86,354,225 47.6 67.0 75.8 106,888,800 88,789,465
MM-Embed(avg. feat.) 52.0 79.0 89.0 103,957,148 86,354,225 43.8 70.1 80.1 106,888,800 88,789,465
jina-clip-v2 57.0 69.3 76.0 92,332,391 83,296,271 45.7 65.1 73.0 94,329,192 86,815,279
3.3 Dataset Statistics
As shown in Table 1, the final dataset of LoVR comprises a total of 467 videos and 40,804 fine-grained clips. Compared to previous benchmarks such as MSR-VTT xu2016msr , MSVD chen2011collecting , and VDC chai2024auroracap , our benchmark offers a significantly larger dataset, longer video durations, and more extensive caption lengths, considering both clip-level and video-level annotations. Annotating with VLM+Human achieves higher scalability than human annotation and better quality than pure VLM-based methods. Furthermore, our benchmark highlights the significance of thematic information, which is neglected by many existing benchmarks. The dataset is divided into a training set and a test set, each containing both long video and clip data. Specifically, to ensure accurate correspondence between clips and captions, we construct the test set by randomly selecting 100 videos to serve as the video portion of the test set. All the clips of these videos form the clip portion of the test set. The remaining 367 videos and all of their associated clips constitute the training set.
4 Experiments
4.1 Experimental Settings
4.1.1 Evaluation Metrics.
We utilize the following two metrics when conducting the quantitative evaluation.
Retrieval Accuracy.
This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results. This metric reflects the recall capability of the retrieval model.
Retrieval Time.
This metric is measured in ms, under two configurations: (1) w/ I/O: End-to-end time including disk read/write and network transmission, reflecting real-world deployment bottlenecks. (2) w/o I/O: Pure computation time excluding I/O operations, such as vector similarity calculation and result ranking, used to assess algorithmic optimization potential.
4.1.2 Baselines
For long video-text retrieval, the encoder plays a central role in modeling visual and text content. Therefore, we select three categories of widely used encoders for evaluating our benchmark. The following are the baselines used:
Image Encoders.
BLIP li2022blip uses an encoder-decoder architecture for vision-language understanding. ResNet50 (CLIP) he2016deep ; radford2021learning is a CNN-based image encoder with skip connections, pre-trained within the CLIP framework. ViT-L/14 (CLIP) dosovitskiy2020image ; radford2021learning is a Vision Transformer trained using the CLIP objective for global image representation.
Video Encoders.
VideoCLIP xu2021videoclip extends CLIP to video-text contrastive learning using temporal aggregation. VideoCLIP-XL is a long-video understanding model designed with high-resolution input to enhance comprehension. VideoCLIP-XL-v2 is an enhanced version of VideoCLIP-XL with better modeling capabilities. wang2024videoclip
Multimodal Embedding Models.
GME-Qwen2-VL-7B-Instruct zhang2024gme is a multilingual multimodal embedding model based on Qwen2-VL. MM-Embedlin2024mm is a universal multimodal retrieval model that integrates multimodal large language capabilities with a modality-aware negative mining strategy. jina-clip-v2 koukounas2024jina extends the CLIP architecture to support multiple languages, high-resolution image inputs, and compact embedding representations using multi-resolution learning.
Table 4: Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
Model Video-to-Text Retrieval Clip-to-Text Retrieval
R@1 R@5 R@10 Time w/ IO Time w/o IO R@1 R@5 R@10 Time w/ IO Time w/o IO
Image Encoder
CLIP(avg. feat.) 19.0 55.0 71.0 21,877,283 3,709,556 27.6 61.7 78.0 22,494,235 3,814,168
BLIP 37.0 60.0 67.0 167,839 18,876 49.4 76.8 85.2 482,842 244,694
ResNet(avg. feat.) 12.0 35.0 60.0 21,096,459 3,016,545 28.9 54.0 66.8 21,691,391 3,101,613
Video Encoder
VideoClip 56.1 79.8 86.3 481,926 6,379 74.1 89.6 93.5 41,867,568 529,587
VideoClip-XL 58.0 80.0 89.0 491,180 6,557 74.8 94.6 97.3 42,835,824 571,837
VideoClip-XL-v2 59.0 81.0 87.0 491,190 6,570 76.0 95.7 98.4 42,836,702 573,026
Multimodal Embedding Model
GME-Qwen2-VL-7B 57.2 83.8 90.2 105,823,692 91,234,289 75.2 93.9 95.1 104,221,391 89,229,158
MM-Embed(avg. feat.) 56.0 84.0 91.0 103,957,148 86,354,225 77.7 94.9 97.6 106,888,800 88,789,465
jina-clip-v2 56.3 83.2 89.4 92,332,391 83,296,271 75.3 92.6 93.4 94,329,192 86,815,279
4.2 Baseline Performance on Video Retrieval and Clip Retrieval
We evaluate a range of baseline methods on the LoVR benchmark across two core tasks: video retrieval and clip retrieval, in both text-to-video/clip and video/clip-to-text directions (see Tables3, and 4). It can be seen that even the top-performing method, MM-Embed (avg. feat.), achieves only 56.0% Recall@1 in video-to-text and 77.7% Recall@1 in clip-to-text retrieval, indicating the difficulty of our benchmark. While models like VideoClip-XL-v2 perform reasonably well on short clips (e.g., 76.0% Recall@1 in clip-to-text), their effectiveness drops notably on full-length videos (59.0% Recall@1). Similarly, image-based encoders such as CLIP show limited generalization to long-form content, with a maximum text-to-video Recall@1 of 69.0%, which is a relatively low score given the complexity of LoVR’s real-world scenes. These findings underscore the unique challenges of the LoVR benchmark, including long temporal dependencies, semantic diversity, and high alignment requirements between modalities, thus providing a meaningful and realistic testbed for advancing robust, scalable long video-text retrieval systems.
4.3 Baseline Performance on Theme Retrieval
As shown in Table 5, we evaluate the performance of baseline models on theme-to-clip and clip-to-theme retrieval tasks. Here, ‘theme’ refers to the thematic information of each clip, focusing on high-level semantic aspects such as narrative themes, emotional tones, and contextual settings.
Considering the ambiguity in thematic information, we expand the evaluation to include Recall@25 and Recall@50. Despite this expansion, the overall performance remains suboptimal, with the highest theme-to-clip Recall@50 score reaching only 77.5%. This suggests that current models still struggle to effectively capture and represent thematic characteristics. Notably, clip-to-theme retrieval performs slightly better than the reverse task. One reason is that the number of distinct themes per clip is relatively limited, which inherently reduces the complexity of the retrieval task.
Table 5: Baseline performance of theme-to-clip and clip-to-theme retrieval. Avg. feat. denotes the feature of the video is the average frame feature.
Model Theme-to-Clip Clip-to-Theme
R@25 R@50 R@25 R@50
CLIP(avg. feat.) 34.3 41.9 43.5 57.2
BLIP 68.1 77.5 61.1 70.5
ResNet 14.9 21.3 - -
ResNet(avg. feat.) 23.9 30.9 42.7 58.0
VideoClip 39.7 52.9 73.2 82.1
VideoClip-XL 45.3 56.1 79.8 86.4
VideoClip-XL-v2 49.3 58.2 80.0 85.6
GME-Qwen2-VL-7B 43.9 49.8 74.6 83.9
MM-Embed 38.2 45.6 - -
MM-Embed(avg. feat.) 47.4 55.4 75.3 85.5
jina-clip-v2 38.3 45.1 73.1 82.4
Figure 5: Case Study of LoVR. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
4.4 Case Study
As shown in Figure 5, two example failure cases are presented. In Query 1, due to the length of the input text, we omit the content of the middle portion. It can be observed that under long-query conditions, the model fails to process the entire context, resulting in the neglect of information at the end of the query. This leads to the omission of critical details — for instance, the key element "pan" in this case is ignored. In Query 2, the model demonstrates limitations in capturing theme and atmosphere information. Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals. These cases highlight the importance of thematic and affective information in video understanding and retrieval tasks.
5 Conclusion
In this work, we present LoVR, a novel benchmark tailored for long video-text retrieval, addressing key limitations in existing datasets such as short video duration, low-quality captions, and insufficient annotation granularity. LoVR features 467 long videos and over 40,000 high-quality, fine-grained clips, supported by a scalable and effective caption generation pipeline that combines VLM-based generation, automatic quality assessment, and iterative refinement. Additionally, we propose a semantic fusion method to produce coherent full-video captions, preserving crucial contextual information. Extensive evaluations using state-of-the-art embedding models demonstrate the challenges posed by LoVR and highlight the limitations of current retrieval methods. We believe that LoVR will serve as a valuable resource for advancing research in multimodal video understanding and retrieval.
Appendix A Details of LoVR Construction
A.1 Implementation Details.
VLM settings
VLMs utilized for caption generation include Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct bai2025qwen2 , and Qwen2-VL-72B-Instruct wang2024qwen2 . The initial temperature is set to 0.9 during caption generation.
Experiment settings
The inter-frame difference scores shown in Figure 3 are calculated by computing the differences between the average pixel values of consecutive frames. To ensure a robust evaluation, we run each quantitative experiment three times and take the average. In the baseline approach, all encoder models first vectorize the videos and the query captions into embeddings. Then, the cosine similarity between each video feature and the query caption feature is computed and used for ranking. The models can be categorized into two groups based on their architecture and processing capabilities:
  • Image encoders and multimodal models that can not process video: These models include CLIP, ResNet, and MM-Embed, which are originally designed for image-level representation and require adaptation to handle video inputs. During the experiments, we set the number of extracted frames used by these models to 10. Two strategies are commonly applied:
    1. Frame-wise maximum similarity: Frames are sampled from each video, and individual frame embeddings are extracted. The cosine similarity between each frame and the query caption is calculated, and the highest score across all frames is taken as the overall similarity between the video and the caption.
    2. Average feature pooling (‘avg. feat.’): Similar to the above, but instead of selecting the maximum similarity, the frame embeddings are averaged to form a single video-level embedding, which is then used to compute similarity with the caption. This strategy is explicitly marked with the suffix ‘avg. feat.’ in Tables 3, 4, and 5. Models without this suffix follow the first strategy.
  • Video encoders and multimodal models that can process video: These models include BLIP, VideoClip, VideoClip-XL, GME, and jina-clip, which are inherently capable of processing temporal information in videos. They directly encode the entire video into a single video-level embedding, based on which the similarity with the text query is computed, without requiring manual aggregation or sampling.
A.2 Computational Costs.
The caption generation process using Qwen2.5-VL-72B-Instruct bai2025qwen2 required approximately 720 GPU hours. Evaluating the generated captions with the EVQAScore liang2024evqascore model added approximately 100 GPU hours to the total computational load. Consequently, the construction of LoVR incurred a total computational cost of roughly 820 GPU hours, based on NVIDIA H800 GPU usage.
Appendix B Extra Experiment Analysis
B.1 Duration of Clips
As shown in the left plot of Figure6, the video duration distribution in our dataset exhibits a balanced coverage of both short and medium-length videos, distinguishing it from existing benchmarks like MSVD chen2011collecting , MSR-VTT xu2016msr and TGIF tgif-cvpr2016 that primarily focus on short clips. Our dataset includes a significant number of medium-length videos (40-120 seconds) that account for approximately 15% of the total collection. This distribution schema enables comprehensive evaluation of video retrieval models, testing their capability to handle both brief visual semantics in short clips and complex temporal relationships in longer videos.
Figure 6: Left: distribution of clip durations in the LoVR dataset. Right: comparison of Recall@K performance among baseline models.
B.2 Comparative Analysis of Recall@K for Baseline Models
The Recall@K curves(right portion of Figure 6)across different models at the clip-level reveal notable performance differences, reflecting variations in model architecture, training strategies, and semantic understanding capabilities. Specifically, the VideoClip-XL series demonstrates the best overall performance, with VideoClip-XL-v2 achieving a Recall@1 of 56.5% and a Recall@20 of 91.5%, significantly outperforming other models. This indicates its stronger ability in fine-grained video understanding. In contrast, early-stage models such as CLIP and ResNet exhibit relatively low accuracy even at small k values, with slower growth trends, highlighting their limitations in capturing complex video semantics.
However, despite some models showing promising results at higher k values, most models still fall short in achieving satisfactory performance on metrics like Recall@1 or Recall@5. This suggests that there is substantial room for improvement, particularly in modeling semantic details and dynamic scene transitions. These results further confirm the high level of challenge posed by our benchmark, which effectively discriminates between model capabilities and provides a meaningful testbed for future research in video-language understanding.
B.3 Impact of Frame Count in Video Embedding on Retrieval Performance
Table 6: Baseline performance of text-to-video and video-to-text retrieval under different frame counts.
Model Frame Count Text-to-Video Retrieval Video-to-Text Retrieval
R@1 R@5 R@10 R@1 R@5 R@10
ResNet 10 32.0 72.0 81.0 12.0 35.0 59.0
20 31.0 72.0 81.0 12.0 35.0 59.0
Clip 10 46.0 78.0 89.0 19.0 55.0 71.0
20 46.0 79.0 89.0 19.0 54.0 71.0
Table 7: Baseline performance of text-to-clip and clip-to-text retrieval under different frame counts.
Model Frame Count Text-to-Clip Retrieval Clip-to-Text Retrieval
R@1 R@5 R@10 R@1 R@5 R@10
ResNet 10 15.1 34.5 44.4 28.8 54.3 66.6
20 15.2 34.9 43.6 29.0 54.0 66.1
Clip 10 25.8 51.6 61.2 27.4 61.9 77.8
20 25.8 51.1 61.4 27.5 61.2 77.7
From Tables 6, 7, and 8, we can observe that the retrieval performance varies under different frame counts, indicating that the number of video frames has an impact on the effectiveness of multimodal retrieval tasks. Although there are slight fluctuations in some metrics, the overall trend suggests that incorporating more frames helps improve the model’s ability to understand and represent video content, thereby enhancing cross-modal matching accuracy. Since videos are composed of multiple frames, a higher frame count generally means more comprehensive visual information. However, using more frames also entails greater computational costs and longer retrieval times, which requires a trade-off based on practical considerations.
At the same time, the results across all tasks reaffirm the high level of challenge posed by the LoVR benchmark dataset. Even when using well-established visual feature extractors such as ResNet, the overall retrieval performance still leaves significant room for improvement, particularly in clip-level retrieval tasks. This highlights the need for stronger cross-modal alignment capabilities in the models.
Table 8: Baseline performance of theme-to-clip and clip-to-theme retrieval under different frame counts.
Model Frame Count Theme-to-Video Retrieval Video-to-Theme Retrieval
R@25 R@50 R@25 R@50
ResNet 10 24.6 31.8 42.9 57.5
20 24.0 31.3 42.7 57.8
Clip 10 34.7 42.1 43.6 57.4
20 35.6 42.4 43.5 57.7
Appendix C Details of Human Quality Assessment
In order to ensure the overall quality and reliability of the video-caption pairs in our dataset, we conducted multiple human evaluation processes. This included: (1) video filter assessment, (2) threshold determination process, (3) human annotation, and (4) final quality assessment.
C.1 High-scene-dynamic Video Filter Assessment
This step is conducted after filtering out the high-scene-dynamic videos (Section 3.2.1). We recruited 10 undergraduate student volunteers with computer science backgrounds to manually review all videos that had passed the filter. The goal was to ensure that only high-dynamic videos were retained in the final dataset. Here, high-dynamic videos are defined as those containing significant scene changes, camera transitions, moving characters or objects, and diverse semantic content.
To enhance the reliability of this filtering stage, each video was independently reviewed by at least two annotators. If both agreed that a video lacked sufficient visual dynamics, it was excluded from the dataset. The entire review process took approximately 2 days, with each participant spending an average of 4 hours on the assigned tasks. A total compensation of ¥3,000 (¥ is the currency symbol for the Chinese yuan (RMB)) was provided to the annotators. As a result of this manual verification, we confirmed and retained 467 high-quality long videos in the final dataset.
C.2 Threshold Determination Process
This step is conducted after scoring the captions generated by VLMs (Section 3.2.2). Each video was assessed by three independent annotators who were tasked with identifying inconsistencies between the captions and the video content. These inconsistencies included textual errors, redundant or missing information, and unaddressed scene transitions. In particular, if a clear scene change occurred in the video (lasting more than 3 seconds) but was not described in the caption, annotators were required to specify its exact duration. Additionally, they evaluated whether the overall theme of the caption significantly deviated from that of the video.
A total of 500 videos were selected for annotation, with each video associated with one clip-level and one video-level caption, resulting in 1,500 annotations in total. Video durations ranged from 3 to 60 seconds. To ensure annotation consistency and quality, six annotators with relevant backgrounds were recruited, each responsible for annotating approximately 250 samples. On average, each annotator spent around 300 minutes on the task. The total project budget was approximately ¥1,200. By analyzing the trends in the manual scores, we were able to preliminarily determine a reasonable threshold for caption quality filtering, which serves as a foundation for subsequent automated evaluation and data curation.
C.3 Human Annotation
This step is conducted to improve the quality of captions getting low scores (Section 3.2.2). To ensure the high quality and consistency of the textual descriptions used in the video-text retrieval tasks, we conducted a detailed annotation process involving a total of 1,171 videos, each ranging from 3 to 60 seconds in duration. The annotation effort was distributed among four annotators, with each annotator reviewing approximately 293 videos. Each annotator received a base payment of ¥250, and those achieving an accuracy rate above 90% on randomly sampled checks (30 per annotator) received an additional bonus of ¥30, leading to a maximum compensation of ¥280 per person.
The annotation guidelines primarily focused on two categories of issues: contextual content errors and topic description discrepancies. Specifically, contextual content errors were further divided into three types:
  • A: Textual expression errors.
  • B: Redundant text.
  • C: Missing key segments, such as unmentioned transitions between clearly distinguishable shots lasting more than 3 seconds.
Topic description discrepancies were categorized as type D, indicating that the overall textual description was clearly inconsistent with the video content.
During annotation, if any error was identified within a sentence, the entire sentence was marked using delimiters ‘’ and ‘’. For type C errors (i.e., missing content), the placeholder ‘ ’ was inserted at the appropriate location where content should be added.
In terms of annotation format, annotators were required to follow a structured labeling protocol when filling out the annotation table:
  • In the “Type” column, the corresponding error types (A, B, C, or D) were recorded. When multiple error types occurred within the same sentence, they were separated by ‘#‘, e.g., A#B#C or B#D.
  • In the “Correction” column, the corrected version of the full sentence was provided. Regardless of how many error types were present in a sentence, only one final corrected version was written, although all relevant error types were listed in the “Type” column.
  • For type B errors (redundant content), if an entire sentence needed to be deleted, the value ‘0‘ was entered in the “Correction” column. If only part of the sentence was redundant, the revised sentence with the unnecessary parts removed was recorded.
  • Annotators were also required to count the number of content-related errors (A+B+C) and topic-level errors (D) for each video and record these statistics in designated fields.
Throughout the annotation workflow, we emphasized consistency in annotation standards to ensure that all corrected texts remained semantically fluent, logically coherent, and accurately reflected the visual content of the videos after replacing the original captions. This annotation mechanism not only improved the overall quality of the dataset but also provided a reliable foundation for subsequent model training and evaluation.
C.4 Final Quality Assessment
As shown in Section 3.2.4, to evaluate the semantic alignment between videos and their corresponding captions in the final dataset (LoVR), we conducted a comprehensive human evaluation involving 25 carefully selected participants. All participants were students with backgrounds in computer science.
Prior to the evaluation, all participants received a detailed instruction manual and completed a short training session using sample videos to standardize their understanding of the scoring criteria. A total of 300 clips and 100 long videos were randomly sampled from the dataset for evaluation. Each sample consisted of a video (or clip) paired with its corresponding caption. Participants were asked to rate the semantic alignment between the visual content and the caption on a 6-point Likert scale (0–5), where:
  • 5: Perfect match — no factual or descriptive errors.
  • 4: Minor issues but overall accurate.
  • 3: Moderate mismatch or ambiguity.
  • 2: Significant inaccuracies.
  • 1: Nearly irrelevant.
  • 0: Completely unrelated.
Each participant was assigned approximately 16 clips and 4 long videos. On average, about 180 seconds were spent per clip, including time for reading instructions and providing ratings. The entire annotation task took roughly one hour per participant. Participants were compensated at a rate of ¥60 per hour.
Figure 7: Prompt used for clip caption generation.
Figure 8: Prompt used for caption aggregation.
Figure 9: Prompt used for sensitive content detection.
Appendix D LLM Prompt
Figure 7 shows the prompt used for clip caption generation, Figure 8 shows the prompt used for caption aggregation, and Figure 8 shows the prompt used for sensitive content detection of captions.
Appendix E Impacts and Safeguards
E.1 Broader Impacts
Our benchmark LoVR does not inherently pose direct negative societal impacts. First, our dataset is sourced from the open-source dataset LongVideoBench wu2024longvideobench , and we strictly adhere to the open-source license specified by LongVideoBench wu2024longvideobench . Furthermore, to mitigate any potential risks, we have taken proactive measures during dataset construction. Specifically, we carefully filtered all source videos to exclude those containing sensitive information(Section C.1). After the final captions are generated, LLM is prompted to detect potential sensitive information in the captions. If any caption is found to contain sensitive content, it is removed to ensure ethical and responsible dataset usage (shown in Figure 9). Although LoVR does not include high-risk models (e.g., image generators or large-scale language models), we still recommend that users follow responsible AI principles when training and deploying models based on LoVR. These principles include, but are not limited to: restricting commercial misuse, conducting fairness assessments, and ensuring transparent and traceable data usage practices.
For the multimodal community, our benchmark not only serves as a comprehensive testbed for video-text retrieval, but the data can also be used as training sets for model development. By introducing long video understanding and retrieval as a key focus area, our work fills a gap in existing benchmarks and provides a valuable reference for future research and advancement in this field.
E.2 Safeguards
To ensure the ethical and responsible release of the LoVR dataset, we have implemented the following safeguards during its construction:
  • All videos are sourced from the open-source benchmark LongVideoBench wu2024longvideobench , and we strictly adhere to its licensing terms.
  • Videos containing sensitive content, such as personally identifiable information, explicit material, or private conversations, are filtered out.
  • A LLM-based sensitive content detector was employed to identify and remove captions that may contain biased, harmful, or inappropriate language.
  • In both the Github and HuggingFace repositories, we recommend that users adopt responsible AI practices when utilizing LoVR, including restricting commercial misuse, performing fairness audits, and ensuring transparent and traceable data usage.
  • On the HuggingFace repository, we implemented a Gated User Access mechanism to ensure that access requests are reviewed and approved, promoting accountable usage of the dataset.
Appendix F Limitations
High computational cost
Our benchmark relies on VLMs during construction, which are inherently resource-intensive in terms of both computation and cost. Moreover, due to the high-dimensional nature of video data, comprising rich spatiotemporal information, the processes of feature extraction, encoding, and storage become particularly demanding when applied to large-scale benchmarks such as LoVR. Additionally, the relatively long video durations and caption lengths in our benchmark further increase the computational burden, introducing additional challenges while also significantly raising resource requirements.
Cross-lingual and cultural gaps
All captions in our benchmark are generated in English, limiting their applicability to non-English scenarios and hindering cross-lingual evaluation. Furthermore, visual expressions, such as gestures, attire, and behavioral patterns, can vary substantially across cultures. However, our caption generation process employs general-purpose models that may fail to capture fine-grained cultural nuances. This limitation affects the benchmark’s ability to reflect culturally specific content, thereby constraining its effectiveness in cross-cultural retrieval settings.

### v2

@@
 LoVR: A Benchmark for Long Video Retrieval in Multimodal Contexts
-Abstract
-Long videos contain a vast amount of information, making video-text retrieval an essential and challenging task in multimodal learning.
-However, existing benchmarks suffer from limited video duration, low-quality captions, and coarse annotation granularity, which hinder the evaluation of advanced video-text retrieval methods.
+Abstract.
+Long videos contain a vast amount of information, making video-text retrieval an essential and challenging task in multimodal learning and web-scale search.
+On today’s Web, where users increasingly expect to locate not only relevant pages but also specific long videos or fine-grained clips, existing benchmarks fall short due to limited video duration, low-quality captions, and coarse annotation granularity.
 To address these limitations, we introduce LoVR, a benchmark specifically designed for long video-text retrieval.
@@
 Our benchmark introduces longer videos, more detailed captions, and a larger-scale dataset, presenting new challenges for video understanding and retrieval.
-Extensive experiments on various advanced embedding models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research.
-We release the code and dataset link at https://github.com/TechNomad-ds/LoVR-benchmark
-Figure 1: Example of video data from LoVR.
-A long video consists of multiple clips, three of which are shown on the left.
-The corresponding caption is displayed on the right, describing each clip’s scene and atmosphere, their inter-linkage, and an overall theme summary.
-Table 1: Comparison of statistics of video-text benchmarks.
-Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator Theme Info.
-MSR-VTT xu2016msr 1,000 15.01 9.41 Human
-DiDeMo anne2017localizing 1,037 53.94 29.11 Human
-MSVD chen2011collecting 670 10.04 7.01 Human
-ActivityNet yu2019activitynet 5,044 36.00 13.48 Human
-DREAM-1K wang2024tarsier 1,000 8.90 59.30 Human
-VDC chai2024auroracap 1,000 28.18 500.91 VLM
-CAREBENCH xu2024fine 1,000 14.35 227.95 Human
+Extensive experiments on various advanced models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research.
+We release the code link at https://anonymous.4open.science/r/LoVR-benchmark-A207/.
+Video-Text Retrieval, Multimodal, Benchmark
+[…]
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
+Figure 1. Example from LoVR. A long video is divided into clips, three shown on the left. The caption on the right describes each clip’s scene and atmosphere, their relations, and the overall summary.
+1. Introduction
+Table 1. Comparison of statistics of video-text benchmarks.
+Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator
+MSR-VTT (Xu et al., 2016) 1,000 15.01 9.41 Human
+DiDeMo (Anne Hendricks et al., 2017) 1,037 53.94 29.11 Human
+MSVD (Chen and Dolan, 2011) 670 10.04 7.01 Human
+ActivityNet (Yu et al., 2019) 5,044 36.00 13.48 Human
+DREAM-1K (Wang et al., 2024c) 1,000 8.90 59.30 Human
+VDC (Chai et al., 2024) 1,000 28.18 500.91 VLM
+CAREBENCH (Xu et al., 2024) 1,000 14.35 227.95 Human
 LoVR
-Clip-level 40,804 17.86 1393.65 VLM + Human $\checkmark$
-Video-level 467 1560.30 106075.39 VLM + Human $\checkmark$
-1 Introduction
-In the data-driven era, with the rapid advancement of digital technology, video has become a crucial medium for information transmission due to its intuitive and richly expressive nature pouyanfar2018multimedia .
-As a core task in multimodal understanding, video-text retrieval, which aims to retrieve semantically relevant video content from large-scale datasets given a textual query, has become increasingly important zhu2023deep .
-Recently, videos have grown significantly in length, making retrieving full long-form videos and fine-grained clips increasingly relevant lin2022eclipse ; luo2024video ; eltahir2025multimodal .
-Consequently, there is an urgent need to benchmark both long video retrieval and clip-level retrieval from long videos to support real-world applications.
-However, current video-text retrieval benchmarks are primarily constrained by short video lengths and low-quality captions. These limitations can be summarized in the following three aspects:
-(1) Limited Video Length Most widely used datasets—such as MSR-VTT xu2016msr , MSVD chen2011collecting , and DiDeMo anne2017localizing contain only short videos, with average durations less than 1 minute.
+Clip-level 40,804 17.86 1393.65 VLM + Human
+Video-level 467 1560.30 106075.39 VLM + Human
+With the explosive growth of online media platforms, video has become a dominant content form on the Web.
+Notably, long-form videos, such as lectures, documentaries, and live streams lasting from minutes to hours, constitute a significant and growing portion of this information ecosystem.
+Effective retrieval of such content is therefore essential for modern web search, where users increasingly expect to locate not just relevant videos but specific moments within them.
+In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (Zhu et al., 2023a).
+However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (Lin et al., 2022; Luo et al., 2024; Eltahir et al., 2025).
+This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(Xu et al., 2016), DiDeMo(Anne Hendricks et al., 2017)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization.
+Consequently, there is an urgent need for a dedicated benchmark to drive progress in long-form video and clip-level retrieval.
+The limitations of existing benchmarks can be summarized in three key aspects:
+(1) Limited Video Length Most widely used datasets—such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), and DiDeMo (Anne Hendricks et al., 2017) contain only short videos, with average durations less than 1 minute.
 These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
 (2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance.
-While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) malik2021automatic ), these methods often suffer from low accuracy and semantic ambiguity.
-Although manual annotation wu2024longvideobench ; song2024moviechat ; yu2019activitynet can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
+While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (Malik et al., 2021)), these methods often suffer from low accuracy and semantic ambiguity.
+Although manual annotation (Wu et al., 2024; Song et al., 2024; Yu et al., 2019) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
 Balancing scalability with annotation quality remains a critical challenge in current research.
-(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos eltahir2025multimodal ; zhu2024video .
+(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (Eltahir et al., 2025; Zhu et al., 2024).
 This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content.
-In contrast, constructing clip-level benchmarks by segmenting long videos introduces a more realistic and challenging setting—models must not only retrieve relevant content but also distinguish fine-grained clips from densely packed temporal contexts.
+In contrast, constructing clip-level benchmarks by segmenting long videos introduces a more realistic and challenging setting that models must not only retrieve relevant content but also distinguish fine-grained clips from densely packed temporal contexts.
 This setting better reflects practical applications and poses a greater challenge for retrieval models.
@@
 To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency.
-Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) zhang2024vision , (2) automatically assessing caption quality via the EVQAScore model liang2024evqascore , and (3) conducting a final round of human verification.
+Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (Zhang et al., 2024a), (2) automatically assessing caption quality via the EVQAScore (Liang et al., 2024), and (3) conducting a final round of human verification.
 This hybrid process effectively ensures both high fidelity and scalability of annotations.
-For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model wang2024qwen2 ; bai2025qwen2 , which demonstrates strong performance on multimodal understanding tasks.
-The integration of EVQAScore liang2024evqascore further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
+For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (Wang et al., 2024a; Bai et al., 2025), which demonstrates strong performance on multimodal understanding tasks.
+The integration of EVQAScore (Liang et al., 2024) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
 By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
 Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios.
-First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content ge2024v2pe ; zhao2023mmicl .
+First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (Ge et al., 2024; Zhao et al., 2023).
 As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
 Second, captions play a vital role in aligning textual and visual modalities.
-In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment chen2025expertized .
-Accurately modeling high-level semantics—such as overarching themes, narrative progression, and emotional tone—remains an open challenge faghihi2024spectrum .
+In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (Chen et al., 2025).
+Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (Faghihi et al., 2024).
 These observations underscore the importance of a dedicated benchmark.
@@
 Our contributions are as follows:
-  1. 1.
+Figure 2.
+Overview of the data construction pipeline in LoVR.
+Step 1 segments long videos into high-dynamic clips using visual change detection and threshold-based filtering.
+Step 2 generates high-quality clip-level captions via iterative VLM captioning and human fallback based on EVQAScore.
+Step 3 constructs long-video captions by clustering clip captions and summarizing them into a full-length description.
+Finally, a human review process is conducted to ensure quality.
+  1. (1)
     We introduce LoVR, a novel benchmark designed for long video-text retrieval.
@@
 Table 1 compares LoVR and other existing benchmarks.
-  2. 2.
+  2. (2)
     To ensure caption quality, we propose a general and cost-effective caption generation pipeline that integrates VLMs with both automated and human validation.
 Compared to fully automatic or fully manual annotation, our approach achieves high-quality captions while significantly reducing human labor by effectively balancing machine efficiency and human oversight.
-  3. 3.
-    We conduct systematic experiments using various embedding models on LoVR, providing insights into effective representation learning and retrieval strategies for long-form video-text understanding.
+  3. (3)
+    We conduct systematic experiments using various models on LoVR, providing insights into effective representation learning and retrieval strategies for long-form video-text understanding.
 Our results reveal that LoVR is more challenging than existing benchmarks, especially in the context of long video retrieval, thus introducing a new challenge and setting a fresh research direction for the community.
-Table 2: Duration distribution of long videos in LoVR. ‘Avg.’ refers to the average duration in seconds, while ‘#Clip’ indicates the average number of clips obtained by segmenting each video.
-Duration (900, 1000] (1000, 2000] (2000, 3000] (3000, 3600]
-Statistics Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip Avg. Count #Clip
-998.7 4 95.5 1293.2 377 88.6 2569.2 61 84.6 3215.6 25 75.1
-2 Related Works
+2. Related Works
 Video-text retrieval benchmarks
-Traditional benchmarks (e.g., MSR-VTT xu2016msr , MSVD chen2011collecting , DiDeMo anne2017localizing , TGIF tgif-cvpr2016 , LSMDC lsmdc ) focus on short video clips with concise captions, limiting their ability to evaluate fine-grained semantic understanding.
-Recent efforts (e.g., ActivityNet yu2019activitynet , HowTo100M miech2019howto100m ) scale up video duration and annotation richness, but HowTo100M relies on noisy ASR-generated captions.
-To improve annotation quality, DREAM-1K wang2024tarsier and VDC chai2024auroracap use human-annotated descriptions, yet their manual annotation process hinders scalable expansion.
-To overcome these challenges, we leverage recent advances in multi-concept VLMs an2024mc and visual prompting lin2025drawandunderstand to introduce a VLM-based initial annotator, enabling scalable and fine-grained long-video retrieval.
+Traditional benchmarks such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), DiDeMo (Anne Hendricks et al., 2017) TGIF (Li et al., 2016) and LSMDC (Rohrbach et al., 2017) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding.
+Recent efforts (Yu et al., 2019) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (Miech et al., 2019) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise.
+To address this limitation, DREAM-1K (Wang et al., 2024c) and VDC (Chai et al., 2024) adopt human-annotated descriptions to improve annotation accuracy.
+However, their heavy reliance on manual annotation constrains scalable dataset expansion, as the labor-intensive process complicates large-scale data acquisition.
 Video-text retrieval methods
@@
 For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations.
-zhu2023deeplearningvideotextretrieval ; an2024mc Feature matching has evolved from global(e.g.
-CLIP4Clip luo2021clip4clipempiricalstudyclip ,VSE++ han2022bicnetlearningefficientspatiotemporal ) to multi-granular approaches (local and individual-level alignment)(e.g.
-CLIP2TV gao2022clip2tvalignmatchdistill ), improving retrieval accuracy.
+(Zhu et al., 2023b) Feature matching has evolved from global(e.g.
+CLIP4Clip (Luo et al., 2021),VSE++ (Han et al., 2022)) to multi-granular approaches (local and individual-level alignment)(e.g.
+CLIP2TV (Gao et al., 2022)), improving retrieval accuracy.
 However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
-3 LoVR Datasets Construction
+3. LoVR Datasets Construction
 The LoVR dataset was constructed through a rigorous process involving raw video selection(Section 3.1) and dataset construction pipeline(Section 3.2).
 Section 3.3 provides an overview of LoVR and highlights its key advantages in comparison to existing benchmarks.
-Figure 2: Overview of the three-stage pipeline in LoVR.
-Step 1 segments long videos into high-dynamic clips using visual change detection and threshold-based filtering.
-Step 2 generates high-quality clip-level captions via iterative VLM captioning and human fallback based on EVQAScore.
-Step 3 constructs long-video captions by clustering clip captions and summarizing them into a full-length description.
-Finally, a human review process is conducted to ensure quality.
-3.1 Video Source
-The LoVR benchmark is constructed based on the LongVideoBench wu2024longvideobench dataset.
-LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 3600 seconds (i.e., 1 hour).
+Figure 3.
+Illustration of Video Dynamics: High vs.
+Low Scene Dynamics.
+Example video frames are shown in the upper part (high-dynamic on the left, low-dynamic on the right), while the lower part presents the computed inter-frame difference curves.
+The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
+3.1. Video Source
+The LoVR benchmark is constructed based on the LongVideoBench (Wu et al., 2024) dataset.
+LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 1 hour.
 The domains of the videos can be broadly categorized into four domains: Life, Movie, Knowledge, and News.
 To focus on long-form video retrieval, we apply a duration filter and retain only those videos exceeding 15 minutes (900 seconds) in length.
-Table 2 summarizes the distribution of long videos across different duration intervals, including the average duration and the number of videos in each group.
-3.2 Dataset Construction Pipeline
+3.2. Dataset Construction Pipeline
 As illustrated in Figure 2, the dataset construction pipeline proposed in this paper consists of three key stages: (1) clip segmentation, (2) caption generation, and (3) full video captioning.
@@
 Following these stages, final human quality assessment is conducted to ensure the reliability and quality of the benchmark.
-3.2.1 Clip Segmentation
+3.2.1. Clip Segmentation
 Clips serve as fundamental semantic units of videos.
@@
 Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval.
-We employ the ContentDetector module from PySceneDetect PySceneDetect , which segments videos based on visual content changes across frames.
+We employ the ContentDetector module from PySceneDetect (PySceneDetect, 2021), which segments videos based on visual content changes across frames.
 This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated.
-We empirically set $\tau=34$, following standard practices in scene detection literature.
 Based on visual scene dynamics, videos can be naturally categorized into two types: those with high scene dynamics and those with low scene dynamics.
 As illustrated in Figure 3, videos such as lectures, interviews, or static photography exhibit minimal visual variation throughout their duration, with scenes remaining largely unchanged across different segments.
-The inclusion of such videos simplifies the retrieval task, as only a small number of frames may suffice to represent the entire content.
+Therefore, we set $\tau=34$, segmenting clips with high-scene dynamics while preserving those with low-scene dynamics.
+The inclusion of low-scene dynamic videos simplifies the retrieval task, as only a small number of frames may suffice to represent the entire content.
 In contrast, videos with high scene dynamics contain rich and diverse visual transitions, requiring more precise matching and understanding, which is precisely the level of challenge our benchmark aims to introduce.
@@
 Ultimately, we curated a final set of 467 high-quality long videos that exhibit rich scene dynamics and semantic content diversity, from which a total of 40,804 clips were extracted.
-Figure 3: Illustration of Video Dynamics: High vs.
-Low Scene Dynamics.
-Example video frames are shown in the upper part (high-dynamic on the left, low-dynamic on the right), while the lower part presents the computed inter-frame difference curves.
-The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
-3.2.2 Caption Generation
+Figure 4.
+The figure shows four panels from left to right: (a) Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$.
+The left y-axis shows the match score distribution, while the right y-axis displays the median content and theme annotation errors.
+A notable reduction in annotation errors occurs when the match score exceeds 0.2.
+(b) Distribution of caption lengths for long videos, with most captions concentrated around 10,000 tokens.
+(c) Distribution of clip durations.
+(d) Distribution of caption lengths for video clips.
+3.2.2. Caption Generation
 We propose a three-stage annotation framework: (1) automatic captioning, (2) quality assessment, and (3) dynamic optimization.
@@
 Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM.
-To ensure high caption quality, we employ Qwen2.5-VL-Instruct bai2025qwen2 , a series of advanced VLMs.
+To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (Bai et al., 2025), a series of advanced VLMs.
 When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style.
-Here, we emphasize the importance of VLMs in describing theme information, such as overall visual style and atmosphere, which is crucial for understanding videos but often overlooked.
-As shown in Table 1, existing benchmark captions lack this type of thematic content.
 Quality Assessment.
-To assess the quality of generated captions, we utilize EVQAScore liang2024evqascore , a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
+To assess the quality of generated captions, we utilize EVQAScore (Liang et al., 2024), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
 For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption.
@@
 This allows us to establish an empirical correlation between the computed matching scores and human-verified relevance.
-As illustrated in the left portion of Figure 4, a significant rise in annotation errors is observed for samples with scores below 0.2.
+As illustrated in Figure 4(a), a significant rise in annotation errors is observed for samples with scores below 0.2.
 Based on this analysis, we set the quality threshold to $\theta=0.2$.
 Captions with scores above $\theta$ are retained as high-quality, while those below are flagged as low-quality, needing refinement or exclusion.
-Figure 4: Left: Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$.
-The score distribution refers to the left y-axis, while the Content Error Median and Theme Error Median refer to the right y-axis.
-A significant drop in annotation errors is observed when the match score exceeds 0.2.
-Middle: Caption length distribution of the long videos.
-Video-level captions are very long, centered around 10,000 tokens.
-Right: Caption length distribution of the clips.
 Dynamic Optimization.
-For captions identified as low-quality, we switch to an alternative VLM to regenerate captions, which are then re-evaluated using the EVQAScore liang2024evqascore .
+For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (Wang et al., 2024a).
+The initial temperature is set to 0.9 during caption generation.
+to regenerate captions, which are then re-evaluated using the EVQAScore (Liang et al., 2024).
 Only those captions achieving a score above the quality threshold $\theta$ are retained.
@@
 This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
-3.2.3 Full Video Captioning
+3.2.3. Full Video Captioning
 Due to the length limitations of VLMs, directly generating captions for entire long videos often leads to incomplete or fragmented descriptions.
@@
 The process is shown as follows:
-$\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$ (1)
+(1) $\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$
 where LLM denotes the rewriting process of the LLM.
@@
 During this process, the clip-level captions $\text{cap}_{c}$ are sequentially aggregated to construct the final video-level caption $\text{cap}_{v}$.
-3.2.4 Final Human Quality Assessment
+Table 2. Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
+Model Text-to-Video Retrieval Text-to-Clip Retrieval
+R@1 R@5 R@10 Time w/ IO Time w/o IO R@1 R@5 R@10 Time w/ IO Time w/o IO
+Image Encoder
+CLIP 23.34 43.90 54.39 2544747 1431709 18.78 37.28 46.31 222270298 125052299
+MetaCLIP-ViT-H-14 26.76 53.10 65.52 1165648 313637 25.26 46.10 55.09 101813266 27394590
+siglip-base-patch16-224 19.06 37.90 52.25 1376746 344674 19.34 36.81 45.26 120251547 30105548
+siglip2-base-patch16-224 10.06 27.41 37.04 1336867 269687 8.35 19.03 25.28 116768361 23555767
+EVA02-CLIP-B-16 26.34 51.61 62.96 7491183 5156938 23.37 44.16 53.54 654315581 450431558
+PE-Core-B16-224 23.34 48.60 58.24 7460584 4967311 19.95 38.42 47.30 651642897 433868639
+Video Encoder
+VideoClip-XL 30.84 54.82 65.74 1074042 29435 47.33 72.66 80.81 93812003 2571057
+VideoClip-XL-v2 29.98 55.67 67.67 1073304 29497 55.34 78.29 84.95 93747515 2576415
+Multimodal Embedding Model
+GME-Qwen2-VL 19.91 41.54 53.32 19242308 18301540 53.53 75.89 82.72 1680714676 1598543574
+$M$-Embed 14.13 31.05 40.69 2030535 1168405 32.32 52.84 60.96 177356626 102054123
+LanguageBind-Video 42.61 66.60 77.94 180170 180170 40.30 65.13 74.26 15742371 15742370
+3.2.4. Final Human Quality Assessment
 To assess the quality of the annotations in LoVR, we conducted a human evaluation study involving 25 participants.
@@
 These findings validate the effectiveness of our annotation pipeline and demonstrate that the captions in LoVR accurately reflect the underlying video content.
-Table 3: Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
-Model Text-to-Video Retrieval Text-to-Clip Retrieval
-R@1 R@5 R@10 Time w/ IO Time w/o IO R@1 R@5 R@10 Time w/ IO Time w/o IO
-Image Encoder
-CLIP 69.0 83.0 89.1 21,877,283 3,709,556 22.4 40.6 51.0 22,494,235 3,814,168
-CLIP(avg. feat.) 46.0 79.0 89.0 21,877,283 3,709,556 25.7 51.6 61.4 22,494,235 3,814,168
-BLIP 39.2 59.0 75.0 174,762 41,746 24.7 47.1 55.7 2,595,565 607,738
-ResNet 36.3 58.0 63.0 21,096,459 3,016,545 15.7 27.4 36.0 21,691,391 3,101,613
-ResNet(avg. feat.) 31.0 72.0 82.0 21,096,459 3,016,545 15.1 33.9 44.3 21,691,391 3,101,613
-Video Encoder
-VideoClip 47.2 74.3 82.1 481,926 6,379 45.7 68.2 81.6 41,867,568 529,587
-VideoClip-XL 53.0 80.0 86.1 491,180 6,557 52.5 75.6 84.4 42,835,824 571,837
-VideoClip-XL-v2 57.0 81.2 87.0 491,190 6,570 56.5 79.2 85.4 42,836,702 573,026
-Multimodal Embedding Model
-GME-Qwen2-VL-7B 61.2 73.0 79.0 105,823,692 91,234,289 51.2 68.7 78.4 104,221,391 89,229,158
-MM-Embed 58.1 71.2 78.0 103,957,148 86,354,225 47.6 67.0 75.8 106,888,800 88,789,465
-MM-Embed(avg. feat.) 52.0 79.0 89.0 103,957,148 86,354,225 43.8 70.1 80.1 106,888,800 88,789,465
-jina-clip-v2 57.0 69.3 76.0 92,332,391 83,296,271 45.7 65.1 73.0 94,329,192 86,815,279
-3.3 Dataset Statistics
-As shown in Table 1, the final dataset of LoVR comprises a total of 467 videos and 40,804 fine-grained clips.
-Compared to previous benchmarks such as MSR-VTT xu2016msr , MSVD chen2011collecting , and VDC chai2024auroracap , our benchmark offers a significantly larger dataset, longer video durations, and more extensive caption lengths, considering both clip-level and video-level annotations.
-Annotating with VLM+Human achieves higher scalability than human annotation and better quality than pure VLM-based methods.
-Furthermore, our benchmark highlights the significance of thematic information, which is neglected by many existing benchmarks.
-The dataset is divided into a training set and a test set, each containing both long video and clip data.
-Specifically, to ensure accurate correspondence between clips and captions, we construct the test set by randomly selecting 100 videos to serve as the video portion of the test set.
-All the clips of these videos form the clip portion of the test set.
-The remaining 367 videos and all of their associated clips constitute the training set.
-4 Experiments
-4.1 Experimental Settings
-4.1.1 Evaluation Metrics.
+3.3. Dataset Statistics
+As shown in Table 1, the final dataset of LoVR consists of 467 videos and 40,804 fine-grained clips.
+Compared to conventional benchmarks such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
+The annotation process employs a "VLM + Human" strategy, which achieves greater scalability than pure human annotation and higher quality than purely VLM-based methods.
+Furthermore, in the prompts designed for VLM annotation, the model is explicitly instructed to identify thematic information of the video, serving as contextual reference for video content.
+The entire dataset is designated as the test set, and all data are used during experimental evaluation.
+Figure 5 illustrates key characteristics of the dataset.
+The figure (b) shows that the video captions in our dataset are notably longer and contain more detailed information.
+The figure (c) presents the duration distribution of clip-level videos.
+As the fundamental units of long videos, these clips cover short to medium lengths, and their combination forms the long video structure.
+This contrasts with existing benchmarks like MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and TGIF (Li et al., 2016), which primarily focus on short video clips.
+A considerable proportion of medium-length videos (40–120 seconds) is included, accounting for approximately 15% of the total collection.
+This distribution allows for a comprehensive evaluation of video retrieval models, testing their ability to comprehend brief visual semantics in short clips as well as complex temporal relationships in longer videos.
+The figure (d) from the left displays the distribution of clip caption lengths.
+It can be observed that the captions are highly detailed, with the majority exceeding 1,000 tokens, offering rich descriptions of the clip content.
+4. Experiments
+4.1. Experimental Settings
+4.1.1. Evaluation Metrics.
 We utilize the following two metrics when conducting the quantitative evaluation.
-Retrieval Accuracy.
-This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results. This metric reflects the recall capability of the retrieval model.
-Retrieval Time.
+Retrieval Accuracy
+This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results and reflects the recall capability of the retrieval model.
+For video retrieval with image-based models, we adopt the average feature pooling strategy, where frame-level embeddings are averaged into a single video-level representation before computing similarity with the caption.
+Retrieval Time
 This metric is measured in ms, under two configurations: (1) w/ I/O: End-to-end time including disk read/write and network transmission, reflecting real-world deployment bottlenecks.
 (2) w/o I/O: Pure computation time excluding I/O operations, such as vector similarity calculation and result ranking, used to assess algorithmic optimization potential.
-4.1.2 Baselines
-For long video-text retrieval, the encoder plays a central role in modeling visual and text content.
-Therefore, we select three categories of widely used encoders for evaluating our benchmark.
-The following are the baselines used:
-Image Encoders.
-BLIP li2022blip uses an encoder-decoder architecture for vision-language understanding.
-ResNet50 (CLIP) he2016deep ; radford2021learning is a CNN-based image encoder with skip connections, pre-trained within the CLIP framework.
-ViT-L/14 (CLIP) dosovitskiy2020image ; radford2021learning is a Vision Transformer trained using the CLIP objective for global image representation.
-Video Encoders.
-VideoCLIP xu2021videoclip extends CLIP to video-text contrastive learning using temporal aggregation.
-VideoCLIP-XL is a long-video understanding model designed with high-resolution input to enhance comprehension.
-VideoCLIP-XL-v2 is an enhanced version of VideoCLIP-XL with better modeling capabilities.
-wang2024videoclip
-Multimodal Embedding Models.
-GME-Qwen2-VL-7B-Instruct zhang2024gme is a multilingual multimodal embedding model based on Qwen2-VL.
-MM-Embedlin2024mm is a universal multimodal retrieval model that integrates multimodal large language capabilities with a modality-aware negative mining strategy.
-jina-clip-v2 koukounas2024jina extends the CLIP architecture to support multiple languages, high-resolution image inputs, and compact embedding representations using multi-resolution learning.
-Table 4: Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
+Table 3. Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
 Model Video-to-Text Retrieval Clip-to-Text Retrieval
@@
 Image Encoder
-CLIP(avg. feat.) 19.0 55.0 71.0 21,877,283 3,709,556 27.6 61.7 78.0 22,494,235 3,814,168
-BLIP 37.0 60.0 67.0 167,839 18,876 49.4 76.8 85.2 482,842 244,694
-ResNet(avg. feat.) 12.0 35.0 60.0 21,096,459 3,016,545 28.9 54.0 66.8 21,691,391 3,101,613
+CLIP 16.70 35.55 47.11 2544747 1431709 18.29 35.99 44.76 222270298 125052299
+MetaCLIP-ViT-H-14 22.27 46.68 58.89 1165648 313637 28.32 51.02 60.52 101813266 27394590
+siglip-base-patch16-224 15.84 36.40 45.61 1376746 344674 21.02 40.23 49.50 120251547 30105548
+siglip2-base-patch16-224 7.49 23.55 31.48 1336867 269687 8.21 18.51 24.62 116768361 23555767
+EVA02-CLIP-B-16 16.27 38.97 48.61 7491183 5156938 23.25 44.05 53.83 654315581 450431558
+PE-Core-B16-224 18.42 35.33 43.68 7460584 4967311 18.43 36.00 44.58 651642897 433868639
 Video Encoder
-VideoClip 56.1 79.8 86.3 481,926 6,379 74.1 89.6 93.5 41,867,568 529,587
-VideoClip-XL 58.0 80.0 89.0 491,180 6,557 74.8 94.6 97.3 42,835,824 571,837
-VideoClip-XL-v2 59.0 81.0 87.0 491,190 6,570 76.0 95.7 98.4 42,836,702 573,026
+VideoClip-XL 29.12 54.39 66.38 1074042 29435 44.18 69.95 78.64 93812003 2571057
+VideoClip-XL-v2 25.05 52.03 64.03 1073304 29497 48.54 73.82 81.75 93747515 2576415
 Multimodal Embedding Model
-GME-Qwen2-VL-7B 57.2 83.8 90.2 105,823,692 91,234,289 75.2 93.9 95.1 104,221,391 89,229,158
-MM-Embed(avg. feat.) 56.0 84.0 91.0 103,957,148 86,354,225 77.7 94.9 97.6 106,888,800 88,789,465
-jina-clip-v2 56.3 83.2 89.4 92,332,391 83,296,271 75.3 92.6 93.4 94,329,192 86,815,279
-4.2 Baseline Performance on Video Retrieval and Clip Retrieval
-We evaluate a range of baseline methods on the LoVR benchmark across two core tasks: video retrieval and clip retrieval, in both text-to-video/clip and video/clip-to-text directions (see Tables3, and 4).
-It can be seen that even the top-performing method, MM-Embed (avg.
-feat.), achieves only 56.0% Recall@1 in video-to-text and 77.7% Recall@1 in clip-to-text retrieval, indicating the difficulty of our benchmark.
-While models like VideoClip-XL-v2 perform reasonably well on short clips (e.g., 76.0% Recall@1 in clip-to-text), their effectiveness drops notably on full-length videos (59.0% Recall@1).
-Similarly, image-based encoders such as CLIP show limited generalization to long-form content, with a maximum text-to-video Recall@1 of 69.0%, which is a relatively low score given the complexity of LoVR’s real-world scenes.
-These findings underscore the unique challenges of the LoVR benchmark, including long temporal dependencies, semantic diversity, and high alignment requirements between modalities, thus providing a meaningful and realistic testbed for advancing robust, scalable long video-text retrieval systems.
-4.3 Baseline Performance on Theme Retrieval
-As shown in Table 5, we evaluate the performance of baseline models on theme-to-clip and clip-to-theme retrieval tasks.
-Here, ‘theme’ refers to the thematic information of each clip, focusing on high-level semantic aspects such as narrative themes, emotional tones, and contextual settings.
-Considering the ambiguity in thematic information, we expand the evaluation to include Recall@25 and Recall@50.
-Despite this expansion, the overall performance remains suboptimal, with the highest theme-to-clip Recall@50 score reaching only 77.5%.
-This suggests that current models still struggle to effectively capture and represent thematic characteristics.
-Notably, clip-to-theme retrieval performs slightly better than the reverse task.
-One reason is that the number of distinct themes per clip is relatively limited, which inherently reduces the complexity of the retrieval task.
-Table 5: Baseline performance of theme-to-clip and clip-to-theme retrieval. Avg. feat. denotes the feature of the video is the average frame feature.
-Model Theme-to-Clip Clip-to-Theme
-R@25 R@50 R@25 R@50
-CLIP(avg. feat.) 34.3 41.9 43.5 57.2
-BLIP 68.1 77.5 61.1 70.5
-ResNet 14.9 21.3 - -
-ResNet(avg. feat.) 23.9 30.9 42.7 58.0
-VideoClip 39.7 52.9 73.2 82.1
-VideoClip-XL 45.3 56.1 79.8 86.4
-VideoClip-XL-v2 49.3 58.2 80.0 85.6
-GME-Qwen2-VL-7B 43.9 49.8 74.6 83.9
-MM-Embed 38.2 45.6 - -
-MM-Embed(avg. feat.) 47.4 55.4 75.3 85.5
-jina-clip-v2 38.3 45.1 73.1 82.4
-Figure 5: Case Study of LoVR. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
-4.4 Case Study
+GME-Qwen2-VL 8.35 20.13 28.91 1680714676 1598543574 42.82 66.03 74.15 19242308 18301540
+$M$-Embed 14.99 37.26 48.82 2030535 1168405 32.81 53.55 61.75 177356626 102054123
+LanguageBind-Video 37.47 60.39 73.23 180170 180170 35.83 61.12 70.84 15742371 15742370
+4.1.2. Baselines
+In order to fully evaluate our benchmark, we select three categories of widely used retrieval models. The following are the baselines used:
+Image Encoders.
+CLIP (Radford et al., 2021) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training.
+MetaCLIP-ViT-H-14 (Xu et al., 2023) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation.
+SigLIP-base-patch16-224 (Zhai et al., 2023) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs.
+SigLIP2-base-patch16-224 (Tschannen et al., 2025) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe.
+EVA02-CLIP-B-16 (Fang et al., 2024) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks.
+PE-Core-B16-224 (Bolya et al., 2025) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
+Video Encoders.
+VideoCLIP-XL (Wang et al., 2024b) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content.
+VideoCLIP-XL-v2 (Wang et al., 2024b) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
+Multimodal Embedding Models.
+GME-Qwen2-VL (Zhang et al., 2024b) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs.
+MM-Embed (Lin et al., 2024) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search.
+LanguageBind-Video (Zhu et al., 2023c) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
+4.2. Baseline Performance on Video Retrieval and Clip Retrieval
+We benchmark a comprehensive suite of models on the LoVR dataset across two fundamental retrieval tasks: text-to-video/clip and video/clip-to-text (see Tables 2 and 3).
+Overall, LanguageBind-Video emerges as the most effective model among all evaluated baselines, achieving the highest performance in both text-to-video (42.6% R@1) and video-to-text (37.5% R@1) retrieval.
+Its clip-level performance is also strong, reaching 40.3% R@1 in text-to-clip and 35.8% R@1 in clip-to-text retrieval.
+In contrast, image-based encoders such as CLIP and MetaCLIP-ViT-H-14 show limited scalability when handling long-form videos.
+For instance, CLIP achieves only 23.3% R@1 in text-to-video retrieval and 16.7% R@1 in video-to-text retrieval, demonstrating poor temporal modeling capacity.
+VideoClip-XL-v2, as a dedicated video encoder, exhibits solid performance in short-clip retrieval (55.3% R@1 in text-to-clip and 48.5% R@1 in clip-to-text), but its performance on full-length video retrieval remains moderate (29.9% R@1 in text-to-video and 25.1% R@1 in video-to-text).
+Among multimodal embedding models, GME-Qwen2-VL and MM-Embed demonstrate good clip-level recall (53.5% and 32.3% R@1 in text-to-clip, respectively), yet both suffer a substantial drop on long videos, indicating that multimodal embeddings trained primarily on images struggle to generalize to extended temporal contexts.
+These results collectively highlight the inherent difficulty of the LoVR benchmark.
+Its long temporal dependencies, rich semantic variation, and stringent cross-modal alignment requirements make it a rigorous and realistic testbed for evaluating scalable long video–text retrieval systems.
+Figure 5. Recall@K performance comparison across baseline models on LoVR clip-level retrieval.
+Table 4. Long video retrieval performance under different frame counts across retrieval tasks.
+Model Frames Text-to-Clip Text-to-Video Clip-to-Text Video-to-Text
+R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
+Clip 10 18.82 37.50 46.29 23.34 43.90 54.39 18.16 35.82 44.64 16.49 35.76 46.90
+20 18.94 37.50 46.49 23.13 43.90 54.39 18.20 35.82 44.75 16.49 35.76 46.68
+30 18.78 37.28 46.31 23.34 43.90 54.39 18.20 35.99 44.76 16.70 35.55 47.11
+VideoCLIP-XL 10 49.06 73.69 81.66 30.62 54.60 64.45 45.36 70.96 79.41 30.84 54.39 66.81
+20 48.31 73.34 81.35 30.84 55.46 67.02 44.66 70.67 78.86 29.76 54.60 66.17
+30 47.33 72.67 80.81 30.84 54.82 65.74 44.18 69.95 78.64 29.12 54.39 66.38
+$M$-Embed 10 32.64 53.21 61.35 14.13 31.69 40.90 32.95 53.69 61.80 14.99 37.47 49.25
+20 32.36 53.04 61.20 14.13 31.69 40.69 32.81 53.71 61.76 14.99 37.47 49.25
+30 32.32 52.84 60.96 14.13 31.05 40.69 32.81 53.55 61.75 14.99 37.36 48.82
+4.3. Comparative Analysis of Recall@K for Baseline Models
+Figure 5 presents the Recall@K performance of representative baselines on the clip-level retrieval task.
+VideoClip-XL achieves the strongest overall performance, with Recall@1 of 47.3% and Recall@20 of 87.5%, demonstrating superior temporal modeling and cross-modal alignment.
+LanguageBind follows closely, reaching 40.3% at Recall@1 and 81.9% at Recall@20, while MM-Embed remains moderate, with Recall@1 of 32.3% and Recall@20 below 69.0%.
+In contrast, CLIP performs weakest, starting at only 18.8% and reaching just 55.9% even when $K=20$.
+Notably, despite steady improvement across all models as $K$ increases, even the best-performing model fails to achieve perfect recall at large $K$.
+This plateau effect underscores the intrinsic difficulty of the LoVR benchmark: it requires precise long-range temporal reasoning and robust semantic grounding to retrieve relevant content among diverse, real-world video scenes.
+Consequently, LoVR serves as a rigorous and discriminative benchmark for advancing the next generation of video–language retrieval systems.
+Figure 6. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
+4.4. Impact of Frame Count in Video Embedding on Retrieval Performance
+As shown in Table 4, varying the number of sampled frames leads to only marginal differences in retrieval accuracy.
+While using more frames provides additional visual information, the improvements are inconsistent and often negligible across tasks.
+For example, CLIP shows virtually no improvement when increasing frames from 10 to 30 across all tasks, with performance remaining nearly flat.
+MM-Embed exhibits similar behavior with minimal variations.
+Even VideoCLIP-XL—though achieving the best overall performance—demonstrates mixed results: while Text-to-Video tasks show slight improvements with more frames, Text-to-Clip and Video-to-Text tasks actually perform best with 10 frames, suggesting that simply adding more frames may introduce noise rather than useful information.These results underscore the inherent difficulty of the LoVR benchmark.
+Despite employing advanced visual feature extractors, the retrieval performance remains far from satisfactory, particularly evident in the large performance gaps between different model architectures.
+The benchmark therefore presents a demanding setting that cannot be solved by naive frame sampling strategies alone, but instead calls for more sophisticated algorithms and models capable of robust long-video understanding and cross-modal alignment.
+4.5. Case Study
 As shown in Figure 6, two example failure cases are presented.
@@
 Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals.
-These cases highlight the importance of thematic and affective information in video understanding and retrieval tasks.
-5 Conclusion
+These cases highlight the importance of accurate long context understanding in long video retrieval tasks.
+5. Conclusion
 In this work, we present LoVR, a novel benchmark tailored for long video-text retrieval, addressing key limitations in existing datasets such as short video duration, low-quality captions, and insufficient annotation granularity.
@@
 We believe that LoVR will serve as a valuable resource for advancing research in multimodal video understanding and retrieval.
+Acknowledgements.
+To Robert, for the bagels and explaining CMYK and color spaces.
 Appendix A Details of LoVR Construction
-A.1 Implementation Details.
-VLM settings
-VLMs utilized for caption generation include Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-7B-Instruct bai2025qwen2 , and Qwen2-VL-72B-Instruct wang2024qwen2 .
-The initial temperature is set to 0.9 during caption generation.
-Experiment settings
+A.1. Implementation Details.
 The inter-frame difference scores shown in Figure 3 are calculated by computing the differences between the average pixel values of consecutive frames.
@@
 Then, the cosine similarity between each video feature and the query caption feature is computed and used for ranking.
-The models can be categorized into two groups based on their architecture and processing capabilities:
-  • Image encoders and multimodal models that can not process video: These models include CLIP, ResNet, and MM-Embed, which are originally designed for image-level representation and require adaptation to handle video inputs.
-During the experiments, we set the number of extracted frames used by these models to 10.
-Two strategies are commonly applied:
-    1.
-Frame-wise maximum similarity: Frames are sampled from each video, and individual frame embeddings are extracted.
-The cosine similarity between each frame and the query caption is calculated, and the highest score across all frames is taken as the overall similarity between the video and the caption.
-    2.
-Average feature pooling (‘avg.
-feat.’): Similar to the above, but instead of selecting the maximum similarity, the frame embeddings are averaged to form a single video-level embedding, which is then used to compute similarity with the caption.
-This strategy is explicitly marked with the suffix ‘avg.
-feat.’ in Tables 3, 4, and 5.
-Models without this suffix follow the first strategy.
-  • Video encoders and multimodal models that can process video: These models include BLIP, VideoClip, VideoClip-XL, GME, and jina-clip, which are inherently capable of processing temporal information in videos.
-They directly encode the entire video into a single video-level embedding, based on which the similarity with the text query is computed, without requiring manual aggregation or sampling.
-A.2 Computational Costs.
-The caption generation process using Qwen2.5-VL-72B-Instruct bai2025qwen2 required approximately 720 GPU hours.
-Evaluating the generated captions with the EVQAScore liang2024evqascore model added approximately 100 GPU hours to the total computational load.
-Consequently, the construction of LoVR incurred a total computational cost of roughly 820 GPU hours, based on NVIDIA H800 GPU usage.
-Appendix B Extra Experiment Analysis
-B.1 Duration of Clips
-As shown in the left plot of Figure6, the video duration distribution in our dataset exhibits a balanced coverage of both short and medium-length videos, distinguishing it from existing benchmarks like MSVD chen2011collecting , MSR-VTT xu2016msr and TGIF tgif-cvpr2016 that primarily focus on short clips.
-Our dataset includes a significant number of medium-length videos (40-120 seconds) that account for approximately 15% of the total collection.
-This distribution schema enables comprehensive evaluation of video retrieval models, testing their capability to handle both brief visual semantics in short clips and complex temporal relationships in longer videos.
-Figure 6: Left: distribution of clip durations in the LoVR dataset. Right: comparison of Recall@K performance among baseline models.
-B.2 Comparative Analysis of Recall@K for Baseline Models
-The Recall@K curves(right portion of Figure 6)across different models at the clip-level reveal notable performance differences, reflecting variations in model architecture, training strategies, and semantic understanding capabilities.
-Specifically, the VideoClip-XL series demonstrates the best overall performance, with VideoClip-XL-v2 achieving a Recall@1 of 56.5% and a Recall@20 of 91.5%, significantly outperforming other models.
-This indicates its stronger ability in fine-grained video understanding.
-In contrast, early-stage models such as CLIP and ResNet exhibit relatively low accuracy even at small k values, with slower growth trends, highlighting their limitations in capturing complex video semantics.
-However, despite some models showing promising results at higher k values, most models still fall short in achieving satisfactory performance on metrics like Recall@1 or Recall@5.
-This suggests that there is substantial room for improvement, particularly in modeling semantic details and dynamic scene transitions.
-These results further confirm the high level of challenge posed by our benchmark, which effectively discriminates between model capabilities and provides a meaningful testbed for future research in video-language understanding.
-B.3 Impact of Frame Count in Video Embedding on Retrieval Performance
-Table 6: Baseline performance of text-to-video and video-to-text retrieval under different frame counts.
-Model Frame Count Text-to-Video Retrieval Video-to-Text Retrieval
-R@1 R@5 R@10 R@1 R@5 R@10
-ResNet 10 32.0 72.0 81.0 12.0 35.0 59.0
-20 31.0 72.0 81.0 12.0 35.0 59.0
-Clip 10 46.0 78.0 89.0 19.0 55.0 71.0
-20 46.0 79.0 89.0 19.0 54.0 71.0
-Table 7: Baseline performance of text-to-clip and clip-to-text retrieval under different frame counts.
-Model Frame Count Text-to-Clip Retrieval Clip-to-Text Retrieval
-R@1 R@5 R@10 R@1 R@5 R@10
-ResNet 10 15.1 34.5 44.4 28.8 54.3 66.6
-20 15.2 34.9 43.6 29.0 54.0 66.1
-Clip 10 25.8 51.6 61.2 27.4 61.9 77.8
-20 25.8 51.1 61.4 27.5 61.2 77.7
-From Tables 6, 7, and 8, we can observe that the retrieval performance varies under different frame counts, indicating that the number of video frames has an impact on the effectiveness of multimodal retrieval tasks.
-Although there are slight fluctuations in some metrics, the overall trend suggests that incorporating more frames helps improve the model’s ability to understand and represent video content, thereby enhancing cross-modal matching accuracy.
-Since videos are composed of multiple frames, a higher frame count generally means more comprehensive visual information.
-However, using more frames also entails greater computational costs and longer retrieval times, which requires a trade-off based on practical considerations.
-At the same time, the results across all tasks reaffirm the high level of challenge posed by the LoVR benchmark dataset.
-Even when using well-established visual feature extractors such as ResNet, the overall retrieval performance still leaves significant room for improvement, particularly in clip-level retrieval tasks.
-This highlights the need for stronger cross-modal alignment capabilities in the models.
-Table 8: Baseline performance of theme-to-clip and clip-to-theme retrieval under different frame counts.
-Model Frame Count Theme-to-Video Retrieval Video-to-Theme Retrieval
-R@25 R@50 R@25 R@50
-ResNet 10 24.6 31.8 42.9 57.5
-20 24.0 31.3 42.7 57.8
-Clip 10 34.7 42.1 43.6 57.4
-20 35.6 42.4 43.5 57.7
-Appendix C Details of Human Quality Assessment
+A.2. Computational Costs.
+The caption generation process using Qwen2.5-VL-72B-Instruct (Bai et al., 2025) requires approximately 720 GPU hours.
+Evaluation of the generated captions with the EVQAScore (Liang et al., 2024) model adds approximately 100 GPU hours to the total computational load.
+Consequently, constructing LoVR incurs a total computational cost of roughly 820 GPU hours on NVIDIA H800 GPUs.
+All experiments are conducted on NVIDIA H800 GPUs.
+Appendix B Details of Human Quality Assessment
 In order to ensure the overall quality and reliability of the video-caption pairs in our dataset, we conducted multiple human evaluation processes.
 This included: (1) video filter assessment, (2) threshold determination process, (3) human annotation, and (4) final quality assessment.
-C.1 High-scene-dynamic Video Filter Assessment
+B.1. High-scene-dynamic Video Filter Assessment
 This step is conducted after filtering out the high-scene-dynamic videos (Section 3.2.1).
@@
 As a result of this manual verification, we confirmed and retained 467 high-quality long videos in the final dataset.
-C.2 Threshold Determination Process
+B.2. Threshold Determination Process
 This step is conducted after scoring the captions generated by VLMs (Section 3.2.2).
@@
 By analyzing the trends in the manual scores, we were able to preliminarily determine a reasonable threshold for caption quality filtering, which serves as a foundation for subsequent automated evaluation and data curation.
-C.3 Human Annotation
+B.3. Human Annotation
 This step is conducted to improve the quality of captions getting low scores (Section 3.2.2).
@@
 This annotation mechanism not only improved the overall quality of the dataset but also provided a reliable foundation for subsequent model training and evaluation.
-C.4 Final Quality Assessment
+B.4. Final Quality Assessment
 As shown in Section 3.2.4, to evaluate the semantic alignment between videos and their corresponding captions in the final dataset (LoVR), we conducted a comprehensive human evaluation involving 25 carefully selected participants.
@@
 Participants were compensated at a rate of ¥60 per hour.
-Figure 7: Prompt used for clip caption generation.
-Figure 8: Prompt used for caption aggregation.
-Figure 9: Prompt used for sensitive content detection.
-Appendix D LLM Prompt
+Figure 7. Prompt used for clip caption generation.
+Figure 8. Prompt used for caption aggregation.
+Figure 9. Prompt used for sensitive content detection.
+Appendix C LLM Prompt
 Figure 7 shows the prompt used for clip caption generation, Figure 8 shows the prompt used for caption aggregation, and Figure 8 shows the prompt used for sensitive content detection of captions.
-Appendix E Impacts and Safeguards
-E.1 Broader Impacts
-Our benchmark LoVR does not inherently pose direct negative societal impacts.
-First, our dataset is sourced from the open-source dataset LongVideoBench wu2024longvideobench , and we strictly adhere to the open-source license specified by LongVideoBench wu2024longvideobench .
-Furthermore, to mitigate any potential risks, we have taken proactive measures during dataset construction.
-Specifically, we carefully filtered all source videos to exclude those containing sensitive information(Section C.1).
-After the final captions are generated, LLM is prompted to detect potential sensitive information in the captions.
-If any caption is found to contain sensitive content, it is removed to ensure ethical and responsible dataset usage (shown in Figure 9).
-Although LoVR does not include high-risk models (e.g., image generators or large-scale language models), we still recommend that users follow responsible AI principles when training and deploying models based on LoVR.
-These principles include, but are not limited to: restricting commercial misuse, conducting fairness assessments, and ensuring transparent and traceable data usage practices.
-For the multimodal community, our benchmark not only serves as a comprehensive testbed for video-text retrieval, but the data can also be used as training sets for model development.
-By introducing long video understanding and retrieval as a key focus area, our work fills a gap in existing benchmarks and provides a valuable reference for future research and advancement in this field.
-E.2 Safeguards
-To ensure the ethical and responsible release of the LoVR dataset, we have implemented the following safeguards during its construction:
-  • All videos are sourced from the open-source benchmark LongVideoBench wu2024longvideobench , and we strictly adhere to its licensing terms.
-  • Videos containing sensitive content, such as personally identifiable information, explicit material, or private conversations, are filtered out.
-  • A LLM-based sensitive content detector was employed to identify and remove captions that may contain biased, harmful, or inappropriate language.
-  • In both the Github and HuggingFace repositories, we recommend that users adopt responsible AI practices when utilizing LoVR, including restricting commercial misuse, performing fairness audits, and ensuring transparent and traceable data usage.
-  • On the HuggingFace repository, we implemented a Gated User Access mechanism to ensure that access requests are reviewed and approved, promoting accountable usage of the dataset.
-Appendix F Limitations
-High computational cost
-Our benchmark relies on VLMs during construction, which are inherently resource-intensive in terms of both computation and cost.
-Moreover, due to the high-dimensional nature of video data, comprising rich spatiotemporal information, the processes of feature extraction, encoding, and storage become particularly demanding when applied to large-scale benchmarks such as LoVR.
-Additionally, the relatively long video durations and caption lengths in our benchmark further increase the computational burden, introducing additional challenges while also significantly raising resource requirements.
-Cross-lingual and cultural gaps
-All captions in our benchmark are generated in English, limiting their applicability to non-English scenarios and hindering cross-lingual evaluation.
-Furthermore, visual expressions, such as gestures, attire, and behavioral patterns, can vary substantially across cultures.
-However, our caption generation process employs general-purpose models that may fail to capture fine-grained cultural nuances.
-This limitation affects the benchmark’s ability to reflect culturally specific content, thereby constraining its effectiveness in cross-cultural retrieval settings.

### v3

@@
 Extensive experiments on various advanced models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research.
-We release the code link at https://anonymous.4open.science/r/LoVR-benchmark-A207/.
+We release the code link at https://github.com/TechNomad-ds/LoVR-benchmark/.
 Video-Text Retrieval, Multimodal, Benchmark
@@
 Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator
-MSR-VTT (Xu et al., 2016) 1,000 15.01 9.41 Human
-DiDeMo (Anne Hendricks et al., 2017) 1,037 53.94 29.11 Human
-MSVD (Chen and Dolan, 2011) 670 10.04 7.01 Human
-ActivityNet (Yu et al., 2019) 5,044 36.00 13.48 Human
-DREAM-1K (Wang et al., 2024c) 1,000 8.90 59.30 Human
-VDC (Chai et al., 2024) 1,000 28.18 500.91 VLM
-CAREBENCH (Xu et al., 2024) 1,000 14.35 227.95 Human
+MSR-VTT (xu2016msr) 1,000 15.01 9.41 Human
+DiDeMo (anne2017localizing) 1,037 53.94 29.11 Human
+MSVD (chen2011collecting) 670 10.04 7.01 Human
+ActivityNet (yu2019activitynet) 5,044 36.00 13.48 Human
+DREAM-1K (wang2024tarsier) 1,000 8.90 59.30 Human
+VDC (chai2024auroracap) 1,000 28.18 500.91 VLM
+CAREBENCH (xu2024fine) 1,000 14.35 227.95 Human
 LoVR
@@
 Effective retrieval of such content is therefore essential for modern web search, where users increasingly expect to locate not just relevant videos but specific moments within them.
-In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (Zhu et al., 2023a).
-However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (Lin et al., 2022; Luo et al., 2024; Eltahir et al., 2025).
-This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(Xu et al., 2016), DiDeMo(Anne Hendricks et al., 2017)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization.
+In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (zhu2023deep).
+However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (lin2022eclipse; luo2024video; eltahir2025multimodal).
+This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(xu2016msr), DiDeMo(anne2017localizing)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization.
 Consequently, there is an urgent need for a dedicated benchmark to drive progress in long-form video and clip-level retrieval.
 The limitations of existing benchmarks can be summarized in three key aspects:
-(1) Limited Video Length Most widely used datasets—such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), and DiDeMo (Anne Hendricks et al., 2017) contain only short videos, with average durations less than 1 minute.
+(1) Limited Video Length Most widely used datasets—such as MSR-VTT (xu2016msr), MSVD (chen2011collecting), and DiDeMo (anne2017localizing) contain only short videos, with average durations less than 1 minute.
 These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
 (2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance.
-While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (Malik et al., 2021)), these methods often suffer from low accuracy and semantic ambiguity.
-Although manual annotation (Wu et al., 2024; Song et al., 2024; Yu et al., 2019) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
+While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (malik2021automatic)), these methods often suffer from low accuracy and semantic ambiguity.
+Although manual annotation (wu2024longvideobench; song2024moviechat; yu2019activitynet) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
 Balancing scalability with annotation quality remains a critical challenge in current research.
-(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (Eltahir et al., 2025; Zhu et al., 2024).
+(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (eltahir2025multimodal; zhu2024video).
 This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content.
@@
 To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency.
-Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (Zhang et al., 2024a), (2) automatically assessing caption quality via the EVQAScore (Liang et al., 2024), and (3) conducting a final round of human verification.
+Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (zhang2024vision), (2) automatically assessing caption quality via the EVQAScore (liang2024evqascore), and (3) conducting a final round of human verification.
 This hybrid process effectively ensures both high fidelity and scalability of annotations.
-For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (Wang et al., 2024a; Bai et al., 2025), which demonstrates strong performance on multimodal understanding tasks.
-The integration of EVQAScore (Liang et al., 2024) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
+For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (wang2024qwen2; bai2025qwen2), which demonstrates strong performance on multimodal understanding tasks.
+The integration of EVQAScore (liang2024evqascore) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
 By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
 Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios.
-First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (Ge et al., 2024; Zhao et al., 2023).
+First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (ge2024v2pe; zhao2023mmicl).
 As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
 Second, captions play a vital role in aligning textual and visual modalities.
-In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (Chen et al., 2025).
-Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (Faghihi et al., 2024).
+In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (chen2025expertized).
+Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (faghihi2024spectrum).
 These observations underscore the importance of a dedicated benchmark.
@@
 Video-text retrieval benchmarks
-Traditional benchmarks such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), DiDeMo (Anne Hendricks et al., 2017) TGIF (Li et al., 2016) and LSMDC (Rohrbach et al., 2017) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding.
-Recent efforts (Yu et al., 2019) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (Miech et al., 2019) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise.
-To address this limitation, DREAM-1K (Wang et al., 2024c) and VDC (Chai et al., 2024) adopt human-annotated descriptions to improve annotation accuracy.
+Traditional benchmarks such as MSR-VTT (xu2016msr), MSVD (chen2011collecting), DiDeMo (anne2017localizing) TGIF (tgif-cvpr2016) and LSMDC (lsmdc) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding.
+Recent efforts (yu2019activitynet) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (miech2019howto100m) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise.
+To address this limitation, DREAM-1K (wang2024tarsier) and VDC (chai2024auroracap) adopt human-annotated descriptions to improve annotation accuracy.
 However, their heavy reliance on manual annotation constrains scalable dataset expansion, as the labor-intensive process complicates large-scale data acquisition.
@@
 For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations.
-(Zhu et al., 2023b) Feature matching has evolved from global(e.g.
-CLIP4Clip (Luo et al., 2021),VSE++ (Han et al., 2022)) to multi-granular approaches (local and individual-level alignment)(e.g.
-CLIP2TV (Gao et al., 2022)), improving retrieval accuracy.
+(zhu2023deeplearningvideotextretrieval) Feature matching has evolved from global(e.g.
+CLIP4Clip (luo2021clip4clipempiricalstudyclip),VSE++ (han2022bicnetlearningefficientspatiotemporal)) to multi-granular approaches (local and individual-level alignment)(e.g.
+CLIP2TV (gao2022clip2tvalignmatchdistill)), improving retrieval accuracy.
 However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
@@
 3.1. Video Source
-The LoVR benchmark is constructed based on the LongVideoBench (Wu et al., 2024) dataset.
+The LoVR benchmark is constructed based on the LongVideoBench (wu2024longvideobench) dataset.
 LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 1 hour.
@@
 Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval.
-We employ the ContentDetector module from PySceneDetect (PySceneDetect, 2021), which segments videos based on visual content changes across frames.
+We employ the ContentDetector module from PySceneDetect (PySceneDetect), which segments videos based on visual content changes across frames.
 This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated.
@@
 Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM.
-To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (Bai et al., 2025), a series of advanced VLMs.
+To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (bai2025qwen2), a series of advanced VLMs.
 When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style.
 Quality Assessment.
-To assess the quality of generated captions, we utilize EVQAScore (Liang et al., 2024), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
+To assess the quality of generated captions, we utilize EVQAScore (liang2024evqascore), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
 For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption.
@@
 Dynamic Optimization.
-For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (Wang et al., 2024a).
+For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (wang2024qwen2).
 The initial temperature is set to 0.9 during caption generation.
-to regenerate captions, which are then re-evaluated using the EVQAScore (Liang et al., 2024).
+to regenerate captions, which are then re-evaluated using the EVQAScore (liang2024evqascore).
 Only those captions achieving a score above the quality threshold $\theta$ are retained.
@@
 As shown in Table 1, the final dataset of LoVR consists of 467 videos and 40,804 fine-grained clips.
-Compared to conventional benchmarks such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
+Compared to conventional benchmarks such as MSR-VTT (xu2016msr) and MSVD (chen2011collecting), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
 The annotation process employs a "VLM + Human" strategy, which achieves greater scalability than pure human annotation and higher quality than purely VLM-based methods.
@@
 As the fundamental units of long videos, these clips cover short to medium lengths, and their combination forms the long video structure.
-This contrasts with existing benchmarks like MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and TGIF (Li et al., 2016), which primarily focus on short video clips.
+This contrasts with existing benchmarks like MSVD (chen2011collecting), MSR-VTT (xu2016msr), and TGIF (tgif-cvpr2016), which primarily focus on short video clips.
 A considerable proportion of medium-length videos (40–120 seconds) is included, accounting for approximately 15% of the total collection.
@@
 Image Encoders.
-CLIP (Radford et al., 2021) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training.
-MetaCLIP-ViT-H-14 (Xu et al., 2023) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation.
-SigLIP-base-patch16-224 (Zhai et al., 2023) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs.
-SigLIP2-base-patch16-224 (Tschannen et al., 2025) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe.
-EVA02-CLIP-B-16 (Fang et al., 2024) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks.
-PE-Core-B16-224 (Bolya et al., 2025) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
+CLIP (radford2021learning) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training.
+MetaCLIP-ViT-H-14 (xu2023demystifying) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation.
+SigLIP-base-patch16-224 (zhai2023sigmoid) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs.
+SigLIP2-base-patch16-224 (tschannen2025siglip) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe.
+EVA02-CLIP-B-16 (fang2024eva) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks.
+PE-Core-B16-224 (bolya2025perception) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
 Video Encoders.
-VideoCLIP-XL (Wang et al., 2024b) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content.
-VideoCLIP-XL-v2 (Wang et al., 2024b) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
+VideoCLIP-XL (wang2024videoclip) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content.
+VideoCLIP-XL-v2 (wang2024videoclip) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
 Multimodal Embedding Models.
-GME-Qwen2-VL (Zhang et al., 2024b) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs.
-MM-Embed (Lin et al., 2024) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search.
-LanguageBind-Video (Zhu et al., 2023c) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
+GME-Qwen2-VL (zhang2024gme) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs.
+MM-Embed (lin2024mm) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search.
+LanguageBind-Video (zhu2023languagebind) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
 4.2. Baseline Performance on Video Retrieval and Clip Retrieval
@@
 A.2. Computational Costs.
-The caption generation process using Qwen2.5-VL-72B-Instruct (Bai et al., 2025) requires approximately 720 GPU hours.
-Evaluation of the generated captions with the EVQAScore (Liang et al., 2024) model adds approximately 100 GPU hours to the total computational load.
+The caption generation process using Qwen2.5-VL-72B-Instruct (bai2025qwen2) requires approximately 720 GPU hours.
+Evaluation of the generated captions with the EVQAScore (liang2024evqascore) model adds approximately 100 GPU hours to the total computational load.
 Consequently, constructing LoVR incurs a total computational cost of roughly 820 GPU hours on NVIDIA H800 GPUs.

### v4

@@
 Extensive experiments on various advanced models demonstrate that LoVR is a challenging benchmark, revealing the limitations of current approaches and providing valuable insights for future research.
-We release the code link at https://github.com/TechNomad-ds/LoVR-benchmark/.
+We release the code link at https://lovrbench.github.io/
 Video-Text Retrieval, Multimodal, Benchmark
 […]
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
-Figure 1. Example from LoVR. A long video is divided into clips, three shown on the left. The caption on the right describes each clip’s scene and atmosphere, their relations, and the overall summary.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
 1. Introduction
@@
 Benchmark #Sample Avg. Len.(s) Avg. Captions Annotator
-MSR-VTT (xu2016msr) 1,000 15.01 9.41 Human
-DiDeMo (anne2017localizing) 1,037 53.94 29.11 Human
-MSVD (chen2011collecting) 670 10.04 7.01 Human
-ActivityNet (yu2019activitynet) 5,044 36.00 13.48 Human
-DREAM-1K (wang2024tarsier) 1,000 8.90 59.30 Human
-VDC (chai2024auroracap) 1,000 28.18 500.91 VLM
-CAREBENCH (xu2024fine) 1,000 14.35 227.95 Human
+MSR-VTT (Xu et al., 2016) 1,000 15.01 9.41 Human
+DiDeMo (Anne Hendricks et al., 2017) 1,037 53.94 29.11 Human
+MSVD (Chen and Dolan, 2011) 670 10.04 7.01 Human
+ActivityNet (Yu et al., 2019) 5,044 36.00 13.48 Human
+DREAM-1K (Wang et al., 2024b) 1,000 8.90 59.30 Human
+VDC (Chai et al., 2024) 1,000 28.18 500.91 VLM
+CAREBENCH (Xu et al., 2024) 1,000 14.35 227.95 Human
 LoVR
@@
 Effective retrieval of such content is therefore essential for modern web search, where users increasingly expect to locate not just relevant videos but specific moments within them.
-In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (zhu2023deep).
-However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (lin2022eclipse; luo2024video; eltahir2025multimodal).
-This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(xu2016msr), DiDeMo(anne2017localizing)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization.
+In this context, video-text retrieval has emerged as a core challenge in multimodal understanding and web-scale search systems (Zhu et al., 2023b).
+Figure 1. Example from LoVR. A long video is divided into clips, three shown on the left. The caption on the right describes each clip’s scene and atmosphere, their relations, and the overall summary.
+However, as videos grow in length and complexity, supporting retrieval at both the full-video and fine-grained clip level poses a pressing challenge for real-world applications (Lin et al., 2022; Luo et al., 2024; Eltahir et al., 2025).
+This challenge is exacerbated by the fact that current video-text retrieval benchmarks (e.g., MSR-VTT(Xu et al., 2016), DiDeMo(Anne Hendricks et al., 2017)) are primarily designed around short video clips and often feature captions that lack the temporal specificity needed for clip-level localization.
 Consequently, there is an urgent need for a dedicated benchmark to drive progress in long-form video and clip-level retrieval.
 The limitations of existing benchmarks can be summarized in three key aspects:
-(1) Limited Video Length Most widely used datasets—such as MSR-VTT (xu2016msr), MSVD (chen2011collecting), and DiDeMo (anne2017localizing) contain only short videos, with average durations less than 1 minute.
+(1) Limited Video Length Most widely used datasets—such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), and DiDeMo (Anne Hendricks et al., 2017) contain only short videos, with average durations less than 1 minute.
 These short durations make it difficult to simulate the complexity of real-world long-video retrieval tasks, limiting the ability of such benchmarks to evaluate models on long-form content effectively.
 (2) Low Video Caption Quality The annotation quality of captions significantly affects benchmark performance.
-While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (malik2021automatic)), these methods often suffer from low accuracy and semantic ambiguity.
-Although manual annotation (wu2024longvideobench; song2024moviechat; yu2019activitynet) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
+While many datasets rely on machine-generated annotations (e.g., using Automatic Speech Recognition(ASR) (Malik et al., 2021)), these methods often suffer from low accuracy and semantic ambiguity.
+Although manual annotation (Wu et al., 2024; Song et al., 2024; Yu et al., 2019) can improve semantic fidelity, it typically lacks fine-grained detail and is not scalable to large-scale datasets.
 Balancing scalability with annotation quality remains a critical challenge in current research.
-(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (eltahir2025multimodal; zhu2024video).
+(3) Lack of Clip Level Benchmarks Most existing benchmarks either focus solely on full-video retrieval or rely on pre-segmented short clips that do not originate from continuous long videos (Eltahir et al., 2025; Zhu et al., 2024).
 This limits their ability to evaluate models’ fine-grained temporal localization and contextual understanding within long-form content.
@@
 This setting better reflects practical applications and poses a greater challenge for retrieval models.
-To address the aforementioned limitations, we introduce LoVR, a new benchmark specifically designed for long video-text retrieval.
-Unlike existing datasets that predominantly consist of short video content, LoVR targets the challenges inherent to long-form videos and aims to support both full-video and fine-grained clip-level retrieval.
-As shown in Figure 1, the caption of the full video not only contains holistic information but also corresponds to fine-grained clip-level details.
-To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency.
-Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (zhang2024vision), (2) automatically assessing caption quality via the EVQAScore (liang2024evqascore), and (3) conducting a final round of human verification.
-This hybrid process effectively ensures both high fidelity and scalability of annotations.
-For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (wang2024qwen2; bai2025qwen2), which demonstrates strong performance on multimodal understanding tasks.
-The integration of EVQAScore (liang2024evqascore) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
-By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
-Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios.
-First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (ge2024v2pe; zhao2023mmicl).
-As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
-Second, captions play a vital role in aligning textual and visual modalities.
-In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (chen2025expertized).
-Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (faghihi2024spectrum).
-These observations underscore the importance of a dedicated benchmark.
-By explicitly targeting long-form content, LoVR surfaces technical bottlenecks in current approaches and provides a valuable foundation for developing more robust long video-text retrieval systems.
-Our contributions are as follows:
 Figure 2.
@@
 Finally, a human review process is conducted to ensure quality.
+To address the aforementioned limitations, we introduce LoVR, a new benchmark specifically designed for long video-text retrieval.
+Unlike existing datasets that predominantly consist of short video content, LoVR targets the challenges inherent to long-form videos and aims to support both full-video and fine-grained clip-level retrieval.
+As shown in Figure 1, the caption of the full video not only contains holistic information but also corresponds to fine-grained clip-level details.
+To overcome the shortage of high-quality annotations at scale, we design a cost-effective and scalable caption generation pipeline that balances annotation quality with efficiency.
+Our pipeline comprises three key stages: (1) generating initial captions using a state-of-the-art Vision-Language Model (VLM) (Zhang et al., 2024a), (2) automatically assessing caption quality via the EVQAScore (Liang et al., 2024), and (3) conducting a final round of human verification.
+This hybrid process effectively ensures both high fidelity and scalability of annotations.
+For caption generation, we adopt the advanced Qwen2.5-VL-Instruct model (Wang et al., 2024c; Bai et al., 2025), which demonstrates strong performance on multimodal understanding tasks.
+The integration of EVQAScore (Liang et al., 2024) further enhances the reliability of our automatic validation step by offering fine-grained quality assessments aligned with human judgments.
+By ensuring annotation quality while maintaining scalability, LoVR offers a more realistic and comprehensive evaluation benchmark that better reflects the demands of real-world long video retrieval scenarios.
+Evaluation on the LoVR benchmark reveals several fundamental challenges faced by current video-text retrieval methods in long video scenarios.
+First, the extended temporal span and structural complexity of long videos make it difficult for models to effectively capture and represent all relevant semantic content (Ge et al., 2024; Zhao et al., 2023).
+As video length increases, both the computational cost of feature extraction and index construction rise significantly, while meaningful semantic signals become sparser.
+Second, captions play a vital role in aligning textual and visual modalities.
+In the context of long videos, captions are typically longer and semantically richer, which increases the difficulty of achieving precise video-text alignment (Chen et al., 2025).
+Accurately modeling high-level semantics, such as overarching themes, narrative progression, and emotional tone, remains an open challenge (Faghihi et al., 2024).
+These observations underscore the importance of a dedicated benchmark.
+By explicitly targeting long-form content, LoVR surfaces technical bottlenecks in current approaches and provides a valuable foundation for developing more robust long video-text retrieval systems.
+Our contributions are as follows:
   1. (1)
@@
 Our results reveal that LoVR is more challenging than existing benchmarks, especially in the context of long video retrieval, thus introducing a new challenge and setting a fresh research direction for the community.
-2. Related Works
-Video-text retrieval benchmarks
-Traditional benchmarks such as MSR-VTT (xu2016msr), MSVD (chen2011collecting), DiDeMo (anne2017localizing) TGIF (tgif-cvpr2016) and LSMDC (lsmdc) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding.
-Recent efforts (yu2019activitynet) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (miech2019howto100m) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise.
-To address this limitation, DREAM-1K (wang2024tarsier) and VDC (chai2024auroracap) adopt human-annotated descriptions to improve annotation accuracy.
-However, their heavy reliance on manual annotation constrains scalable dataset expansion, as the labor-intensive process complicates large-scale data acquisition.
-Video-text retrieval methods
-Recent advances have been driven by deep learning techniques.
-For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations.
-(zhu2023deeplearningvideotextretrieval) Feature matching has evolved from global(e.g.
-CLIP4Clip (luo2021clip4clipempiricalstudyclip),VSE++ (han2022bicnetlearningefficientspatiotemporal)) to multi-granular approaches (local and individual-level alignment)(e.g.
-CLIP2TV (gao2022clip2tvalignmatchdistill)), improving retrieval accuracy.
-However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
-3. LoVR Datasets Construction
-The LoVR dataset was constructed through a rigorous process involving raw video selection(Section 3.1) and dataset construction pipeline(Section 3.2).
-Section 3.3 provides an overview of LoVR and highlights its key advantages in comparison to existing benchmarks.
 Figure 3.
@@
 The high-dynamic video exhibits large frame-to-frame variations, whereas the low-dynamic video remains visually stable.
+2. Related Works
+Video-text retrieval benchmarks
+Traditional benchmarks such as MSR-VTT (Xu et al., 2016), MSVD (Chen and Dolan, 2011), DiDeMo (Anne Hendricks et al., 2017) TGIF (Li et al., 2016) and LSMDC (Rohrbach et al., 2017) focus on short video clips paired with concise captions, which limits their capacity to evaluate fine-grained semantic understanding.
+Recent efforts (Yu et al., 2019) aim to scale both the temporal extent of videos and the descriptive richness of annotations, among which HowTo100M (Miech et al., 2019) provides a massive-scale dataset of long-duration tutorial videos, but relies on ASR-generated captions, using only audio to generate labels introduces significant label noise.
+To address this limitation, DREAM-1K (Wang et al., 2024b) and VDC (Chai et al., 2024) adopt human-annotated descriptions to improve annotation accuracy.
+However, their heavy reliance on manual annotation constrains scalable dataset expansion, as the labor-intensive process complicates large-scale data acquisition.
+Video-text retrieval methods
+Recent advances have been driven by deep learning techniques.
+For feature extraction, Transformer-based architectures like ViT and CLIP now dominate, effectively capturing spatiotemporal video features and multimodal data (e.g., audio, OCR), while pretrained language models (e.g., BERT) provide superior text representations.
+(Zhu et al., 2023c) Feature matching has evolved from global(e.g.
+CLIP4Clip (Luo et al., 2021),VSE++ (Han et al., 2022)) to multi-granular approaches (local and individual-level alignment)(e.g.
+CLIP2TV (Gao et al., 2022)), improving retrieval accuracy.
+However, challenges remain in cross-modal alignment and computational efficiency, pointing to future research directions in model optimization and weakly supervised learning.
+3. LoVR Datasets Construction
+The LoVR dataset was constructed through a rigorous process involving raw video selection(Section 3.1) and dataset construction pipeline(Section 3.2).
+Section 3.3 provides an overview of LoVR and highlights its key advantages in comparison to existing benchmarks.
 3.1. Video Source
-The LoVR benchmark is constructed based on the LongVideoBench (wu2024longvideobench) dataset.
+The LoVR benchmark is constructed based on the LongVideoBench (Wu et al., 2024) dataset.
 LongVideoBench contains a diverse collection of videos with durations ranging from 8 seconds to 1 hour.
@@
 Following these stages, final human quality assessment is conducted to ensure the reliability and quality of the benchmark.
+Figure 4.
+The figure shows four panels from left to right: (a) Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$.
+The left y-axis shows the match score distribution, while the right y-axis displays the median content and theme annotation errors.
+A notable reduction in annotation errors occurs when the match score exceeds 0.2.
+(b) Distribution of caption lengths for long videos, with most captions concentrated around 10,000 tokens.
+(c) Distribution of clip durations.
+(d) Distribution of caption lengths for video clips.
 3.2.1. Clip Segmentation
@@
 Segmenting long videos into shorter clips facilitates more fine-grained content analysis and retrieval.
-We employ the ContentDetector module from PySceneDetect (PySceneDetect), which segments videos based on visual content changes across frames.
+We employ the ContentDetector module from PySceneDetect (PySceneDetect, 2021), which segments videos based on visual content changes across frames.
 This approach relies on a threshold $\tau$ to detect scene changes: when the inter-frame difference surpasses $\tau$, a cut is generated.
@@
 Ultimately, we curated a final set of 467 high-quality long videos that exhibit rich scene dynamics and semantic content diversity, from which a total of 40,804 clips were extracted.
-Figure 4.
-The figure shows four panels from left to right: (a) Distribution of match scores and annotation errors on the sampled subset $K^{\prime}$.
-The left y-axis shows the match score distribution, while the right y-axis displays the median content and theme annotation errors.
-A notable reduction in annotation errors occurs when the match score exceeds 0.2.
-(b) Distribution of caption lengths for long videos, with most captions concentrated around 10,000 tokens.
-(c) Distribution of clip durations.
-(d) Distribution of caption lengths for video clips.
 3.2.2. Caption Generation
@@
 Specifically, for each clip $c_{i}\in v$, its caption is generated as $\text{cap}_{i}=\text{VLM}(c_{i})$ where VLM denotes the process of caption generation using VLM.
-To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (bai2025qwen2), a series of advanced VLMs.
+To ensure high caption quality, we employ Qwen2.5-VL-7B-Instruct (Bai et al., 2025), a series of advanced VLMs.
 When generating captions, the VLMs are prompted to describe key aspects of the given clip, including events, scenes, character actions, and overall visual style.
 Quality Assessment.
-To assess the quality of generated captions, we utilize EVQAScore (liang2024evqascore), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
+To assess the quality of generated captions, we utilize EVQAScore (Liang et al., 2024), a fine-grained method designed to measure the semantic alignment between video clips and their corresponding captions.
 For each clip-caption pair $(c_{i},\text{cap}_{i})$, we compute a matching score $s_{i}=\text{EVQAScore}(c_{i},\text{cap}_{i})$, which quantifies the semantic relevance between the visual content and the generated caption.
@@
 Captions with scores above $\theta$ are retained as high-quality, while those below are flagged as low-quality, needing refinement or exclusion.
-Dynamic Optimization.
-For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (wang2024qwen2).
-The initial temperature is set to 0.9 during caption generation.
-to regenerate captions, which are then re-evaluated using the EVQAScore (liang2024evqascore).
-Only those captions achieving a score above the quality threshold $\theta$ are retained.
-This strategy leverages the diversity among VLMs—arising from differences in training data and architectural design—to improve caption quality for challenging clips.
-It is only when the number of failed attempts exceeds three that we resort to manual annotation to generate the final captions.
-This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
-3.2.3. Full Video Captioning
-Due to the length limitations of VLMs, directly generating captions for entire long videos often leads to incomplete or fragmented descriptions.
-To address this issue, we sequentially merge adjacent captions by aligning and fusing semantically related segments rather than naively concatenating all clip-level captions or processing them in an LLM beyond its sequence limit.
-The process is shown as follows:
-(1) $\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$
-where LLM denotes the rewriting process of the LLM.
-In our strategy, we connect the last paragraph of one clip with the first paragraph of the subsequent clip through a language model.
-During this process, the clip-level captions $\text{cap}_{c}$ are sequentially aggregated to construct the final video-level caption $\text{cap}_{v}$.
 Table 2. Baseline performance of text-to-video and text-to-clip retrieval on the LoVR benchmark.
@@
 LanguageBind-Video 42.61 66.60 77.94 180170 180170 40.30 65.13 74.26 15742371 15742370
+Dynamic Optimization.
+For captions identified as low-quality, we switch to alternative VLMs including Qwen2.5-VL-72B-Instruct, and Qwen2-VL-72B-Instruct (Wang et al., 2024c).
+The initial temperature is set to 0.9 during caption generation.
+to regenerate captions, which are then re-evaluated using the EVQAScore (Liang et al., 2024).
+Only those captions achieving a score above the quality threshold $\theta$ are retained.
+This strategy leverages the diversity among VLMs—arising from differences in training data and architectural design—to improve caption quality for challenging clips.
+It is only when the number of failed attempts exceeds three that we resort to manual annotation to generate the final captions.
+This hybrid pipeline reduces the need for human intervention, thereby minimizing annotation cost while ensuring high-quality captioning at scale.
+3.2.3. Full Video Captioning
+Due to the length limitations of VLMs, directly generating captions for entire long videos often leads to incomplete or fragmented descriptions.
+To address this issue, we sequentially merge adjacent captions by aligning and fusing semantically related segments rather than naively concatenating all clip-level captions or processing them in an LLM beyond its sequence limit.
+The process is shown as follows:
+(1) $\text{cap}_{\text{merged}}=p^{\text{rest}}_{\text{c}_{1}}+\text{LLM}\left(p^{\text{tail}}_{\text{c}_{1}}\parallel p^{\text{head}}_{\text{c}_{2}}\right)+p^{\text{rest}}_{\text{c}_{2}}$
+where LLM denotes the rewriting process of the LLM.
+In our strategy, we connect the last paragraph of one clip with the first paragraph of the subsequent clip through a language model.
+During this process, the clip-level captions $\text{cap}_{c}$ are sequentially aggregated to construct the final video-level caption $\text{cap}_{v}$.
 3.2.4. Final Human Quality Assessment
@@
 As shown in Table 1, the final dataset of LoVR consists of 467 videos and 40,804 fine-grained clips.
-Compared to conventional benchmarks such as MSR-VTT (xu2016msr) and MSVD (chen2011collecting), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
+Compared to conventional benchmarks such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), our benchmark provides a larger scale of data, longer video durations, and more comprehensive captions at both the clip level and video level.
 The annotation process employs a "VLM + Human" strategy, which achieves greater scalability than pure human annotation and higher quality than purely VLM-based methods.
@@
 As the fundamental units of long videos, these clips cover short to medium lengths, and their combination forms the long video structure.
-This contrasts with existing benchmarks like MSVD (chen2011collecting), MSR-VTT (xu2016msr), and TGIF (tgif-cvpr2016), which primarily focus on short video clips.
+This contrasts with existing benchmarks like MSVD (Chen and Dolan, 2011), MSR-VTT (Xu et al., 2016), and TGIF (Li et al., 2016), which primarily focus on short video clips.
 A considerable proportion of medium-length videos (40–120 seconds) is included, accounting for approximately 15% of the total collection.
@@
 It can be observed that the captions are highly detailed, with the majority exceeding 1,000 tokens, offering rich descriptions of the clip content.
-4. Experiments
-4.1. Experimental Settings
-4.1.1. Evaluation Metrics.
-We utilize the following two metrics when conducting the quantitative evaluation.
-Retrieval Accuracy
-This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results and reflects the recall capability of the retrieval model.
-For video retrieval with image-based models, we adopt the average feature pooling strategy, where frame-level embeddings are averaged into a single video-level representation before computing similarity with the caption.
-Retrieval Time
-This metric is measured in ms, under two configurations: (1) w/ I/O: End-to-end time including disk read/write and network transmission, reflecting real-world deployment bottlenecks.
-(2) w/o I/O: Pure computation time excluding I/O operations, such as vector similarity calculation and result ranking, used to assess algorithmic optimization potential.
 Table 3. Baseline performance of video-to-text and clip-to-text retrieval on the LoVR benchmark.
@@
 LanguageBind-Video 37.47 60.39 73.23 180170 180170 35.83 61.12 70.84 15742371 15742370
+4. Experiments
+4.1. Experimental Settings
+4.1.1. Evaluation Metrics.
+We utilize the following two metrics when conducting the quantitative evaluation.
+Retrieval Accuracy
+This metric is measured using Recall@K, which quantifies the proportion of true relevant items among the top-K retrieved results and reflects the recall capability of the retrieval model.
+For video retrieval with image-based models, we adopt the average feature pooling strategy, where frame-level embeddings are averaged into a single video-level representation before computing similarity with the caption.
+Retrieval Time
+This metric is measured in ms, under two configurations: (1) w/ I/O: End-to-end time including disk read/write and network transmission, reflecting real-world deployment bottlenecks.
+(2) w/o I/O: Pure computation time excluding I/O operations, such as vector similarity calculation and result ranking, used to assess algorithmic optimization potential.
 4.1.2. Baselines
@@
 Image Encoders.
-CLIP (radford2021learning) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training.
-MetaCLIP-ViT-H-14 (xu2023demystifying) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation.
-SigLIP-base-patch16-224 (zhai2023sigmoid) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs.
-SigLIP2-base-patch16-224 (tschannen2025siglip) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe.
-EVA02-CLIP-B-16 (fang2024eva) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks.
-PE-Core-B16-224 (bolya2025perception) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
+CLIP (Radford et al., 2021) is a foundational vision-language model that learns a shared embedding space for images and text through contrastive pre-training.
+MetaCLIP-ViT-H-14 (Xu et al., 2023) is a large-scale CLIP variant trained on a vast dataset, utilizing a Vision Transformer (ViT-H/14) backbone for enhanced visual representation.
+SigLIP-base-patch16-224 (Zhai et al., 2023) improves upon CLIP by employing a sigmoid-based loss function that operates on individual image-text pairs.
+SigLIP2-base-patch16-224 (Tschannen et al., 2025) is an advanced iteration of SigLIP that integrates multiple pretraining techniques into a unified recipe.
+EVA02-CLIP-B-16 (Fang et al., 2024) leverages the robustly optimized EVA-02 Vision Transformer as its image encoder, offering strong performance in vision-language tasks.
+PE-Core-B16-224 (Bolya et al., 2025) is a competitive perception encoder trained with a robust vision-language learning objective, designed for high performance in image and video understanding tasks.
 Video Encoders.
-VideoCLIP-XL (wang2024videoclip) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content.
-VideoCLIP-XL-v2 (wang2024videoclip) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
+VideoCLIP-XL (Wang et al., 2024a) is specifically designed for long video understanding, leveraging a large-scale dataset of video-long description pairs to enhance its comprehension of extended temporal content.
+VideoCLIP-XL-v2 (Wang et al., 2024a) is an optimized successor to VideoCLIP-XL, trained with additional data and refined techniques to further boost its capabilities in long-form video-text retrieval.
 Multimodal Embedding Models.
-GME-Qwen2-VL (zhang2024gme) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs.
-MM-Embed (lin2024mm) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search.
-LanguageBind-Video (zhu2023languagebind) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
-4.2. Baseline Performance on Video Retrieval and Clip Retrieval
-We benchmark a comprehensive suite of models on the LoVR dataset across two fundamental retrieval tasks: text-to-video/clip and video/clip-to-text (see Tables 2 and 3).
-Overall, LanguageBind-Video emerges as the most effective model among all evaluated baselines, achieving the highest performance in both text-to-video (42.6% R@1) and video-to-text (37.5% R@1) retrieval.
-Its clip-level performance is also strong, reaching 40.3% R@1 in text-to-clip and 35.8% R@1 in clip-to-text retrieval.
-In contrast, image-based encoders such as CLIP and MetaCLIP-ViT-H-14 show limited scalability when handling long-form videos.
-For instance, CLIP achieves only 23.3% R@1 in text-to-video retrieval and 16.7% R@1 in video-to-text retrieval, demonstrating poor temporal modeling capacity.
-VideoClip-XL-v2, as a dedicated video encoder, exhibits solid performance in short-clip retrieval (55.3% R@1 in text-to-clip and 48.5% R@1 in clip-to-text), but its performance on full-length video retrieval remains moderate (29.9% R@1 in text-to-video and 25.1% R@1 in video-to-text).
-Among multimodal embedding models, GME-Qwen2-VL and MM-Embed demonstrate good clip-level recall (53.5% and 32.3% R@1 in text-to-clip, respectively), yet both suffer a substantial drop on long videos, indicating that multimodal embeddings trained primarily on images struggle to generalize to extended temporal contexts.
-These results collectively highlight the inherent difficulty of the LoVR benchmark.
-Its long temporal dependencies, rich semantic variation, and stringent cross-modal alignment requirements make it a rigorous and realistic testbed for evaluating scalable long video–text retrieval systems.
+GME-Qwen2-VL (Zhang et al., 2024b) is a unified multimodal embedding model built upon the Qwen2-VL multimodal large language model, capable of generating joint embeddings for text, images, and image-text pairs.
+MM-Embed (Lin et al., 2024) is a universal multimodal retrieval model based on a multimodal large language model (MLLM) with a bi-encoder architecture, designed for efficient and broad-spectrum cross-modal search.
+LanguageBind-Video (Zhu et al., 2023a) adopts a language-centric approach to align video features with other modalities in a shared semantic space.
 Figure 5. Recall@K performance comparison across baseline models on LoVR clip-level retrieval.
@@
 30 32.32 52.84 60.96 14.13 31.05 40.69 32.81 53.55 61.75 14.99 37.36 48.82
+4.2. Baseline Performance on Video Retrieval and Clip Retrieval
+We benchmark a comprehensive suite of models on the LoVR dataset across two fundamental retrieval tasks: text-to-video/clip and video/clip-to-text (see Tables 2 and 3).
+Overall, LanguageBind-Video emerges as the most effective model among all evaluated baselines, achieving the highest performance in both text-to-video (42.6% R@1) and video-to-text (37.5% R@1) retrieval.
+Its clip-level performance is also strong, reaching 40.3% R@1 in text-to-clip and 35.8% R@1 in clip-to-text retrieval.
+In contrast, image-based encoders such as CLIP and MetaCLIP-ViT-H-14 show limited scalability when handling long-form videos.
+For instance, CLIP achieves only 23.3% R@1 in text-to-video retrieval and 16.7% R@1 in video-to-text retrieval, demonstrating poor temporal modeling capacity.
+VideoClip-XL-v2, as a dedicated video encoder, exhibits solid performance in short-clip retrieval (55.3% R@1 in text-to-clip and 48.5% R@1 in clip-to-text), but its performance on full-length video retrieval remains moderate (29.9% R@1 in text-to-video and 25.1% R@1 in video-to-text).
+Among multimodal embedding models, GME-Qwen2-VL and MM-Embed demonstrate good clip-level recall (53.5% and 32.3% R@1 in text-to-clip, respectively), yet both suffer a substantial drop on long videos, indicating that multimodal embeddings trained primarily on images struggle to generalize to extended temporal contexts.
+These results collectively highlight the inherent difficulty of the LoVR benchmark.
+Its long temporal dependencies, rich semantic variation, and stringent cross-modal alignment requirements make it a rigorous and realistic testbed for evaluating scalable long video–text retrieval systems.
 4.3. Comparative Analysis of Recall@K for Baseline Models
@@
 Consequently, LoVR serves as a rigorous and discriminative benchmark for advancing the next generation of video–language retrieval systems.
-Figure 6. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
 4.4. Impact of Frame Count in Video Embedding on Retrieval Performance
@@
 The benchmark therefore presents a demanding setting that cannot be solved by naive frame sampling strategies alone, but instead calls for more sophisticated algorithms and models capable of robust long-video understanding and cross-modal alignment.
-4.5. Case Study
-As shown in Figure 6, two example failure cases are presented.
-In Query 1, due to the length of the input text, we omit the content of the middle portion.
-It can be observed that under long-query conditions, the model fails to process the entire context, resulting in the neglect of information at the end of the query.
-This leads to the omission of critical details — for instance, the key element "pan" in this case is ignored.
-In Query 2, the model demonstrates limitations in capturing theme and atmosphere information.
-Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals.
-These cases highlight the importance of accurate long context understanding in long video retrieval tasks.
 5. Conclusion
@@
 We believe that LoVR will serve as a valuable resource for advancing research in multimodal video understanding and retrieval.
-Acknowledgements.
-To Robert, for the bagels and explaining CMYK and color spaces.
 Appendix A Details of LoVR Construction
@@
 A.2. Computational Costs.
-The caption generation process using Qwen2.5-VL-72B-Instruct (bai2025qwen2) requires approximately 720 GPU hours.
-Evaluation of the generated captions with the EVQAScore (liang2024evqascore) model adds approximately 100 GPU hours to the total computational load.
+The caption generation process using Qwen2.5-VL-72B-Instruct (Bai et al., 2025) requires approximately 720 GPU hours.
+Evaluation of the generated captions with the EVQAScore (Liang et al., 2024) model adds approximately 100 GPU hours to the total computational load.
 Consequently, constructing LoVR incurs a total computational cost of roughly 820 GPU hours on NVIDIA H800 GPUs.
@@
 This annotation mechanism not only improved the overall quality of the dataset but also provided a reliable foundation for subsequent model training and evaluation.
+Figure 6. Prompt used for clip caption generation.
+Figure 7. Prompt used for caption aggregation.
+Figure 8. Prompt used for sensitive content detection.
 B.4. Final Quality Assessment
@@
 Participants were compensated at a rate of ¥60 per hour.
-Figure 7. Prompt used for clip caption generation.
-Figure 8. Prompt used for caption aggregation.
-Figure 9. Prompt used for sensitive content detection.
 Appendix C LLM Prompt
 Figure 6 shows the prompt used for clip caption generation, Figure 7 shows the prompt used for caption aggregation, and Figure 7 shows the prompt used for sensitive content detection of captions.
+Figure 9. Two illustrative retrieval cases are provided, where the key retrieval targets are highlighted in red.
+Appendix D Case Study
+As shown in Figure 9, two example failure cases are presented.
+In Query 1, due to the length of the input text, we omit the content of the middle portion.
+It can be observed that under long-query conditions, the model fails to process the entire context, resulting in the neglect of information at the end of the query.
+This leads to the omission of critical details — for instance, the key element "pan" in this case is ignored.
+In Query 2, the model demonstrates limitations in capturing theme and atmosphere information.
+Although four videos related to urban city scenes are retrieved, only the first one accurately reflects the required ‘solemn and sorrowful atmosphere.’ The remaining results fail to incorporate this essential emotional aspect, leading to incorrect retrievals.
+These cases highlight the importance of accurate long context understanding in long video retrieval tasks.